### v1

# Metric and Geometric Spanners that are Resilient to Degree-Bounded Edge Faults

Abstract
Let $H$ be an edge-weighted graph, and let $G$ be a subgraph of $H$. We say that $G$ is an $f$-fault-tolerant $t$-spanner for $H$, if the following is true for any subset $F$ of at most $f$ edges of $G$: For any two vertices $p$ and $q$, the shortest-path distance between $p$ and $q$ in the graph $G\setminus F$ is at most $t$ times the shortest-path distance between $p$ and $q$ in the graph $H\setminus F$. The value of $t$ is called the stretch factor. Numerous results are known for constructing such spanners, both for general graphs $H$ and for complete graphs $H$ whose vertex set is a set of points in $\mathbb{R}^{d}$ or in a metric space of bounded doubling dimension.
Recently, Bodwin, Haeupler, and Parter generalized this notion to the case when $F$ can be any set of edges in $G$, as long as the maximum degree of $F$ is at most $f$. For example, if $f=1$, then $F$ can be a perfect matching in $G$. They gave constructions for general graphs $H$.
We first consider the case when $H$ is a complete graph whose vertex set is an arbitrary metric space. We show that if this metric space contains a $t$-spanner with $m$ edges, then it also contains a graph $G$ with at most $(4f-1)m$ edges, that is resilient to edge faults of maximum degree $f$ and has stretch factor $O(ft)$. We also give an example for which our construction has stretch factor $\Omega(ft)$. If $f=1$, the graph $G$ has at most $3m$ edges and its stretch factor is at most $3t$. For this case, we give an example for which $G$ has stretch factor arbitrarily close to $3t$.
Next, we consider the case when $H$ is a complete graph whose vertex set is a metric space that admits a well-separated pair decomposition. We show that, if the metric space has such a decomposition of size $m$, then it contains a graph with at most $(2f+1)^{2}m$ edges, that is resilient to edge faults of maximum degree $f$ and has stretch factor at most $1+\varepsilon$, for any given $\varepsilon>0$. For example, if the vertex set is a set of $n$ points in $\mathbb{R}^{d}$ ($d$ being a constant) or a set of $n$ points in a metric space of bounded doubling dimension, then the spanner has $O(f^{2}n)$ edges.
Finally, for the case when $H$ is a complete graph on $n$ points in $\mathbb{R}^{d}$, we show how natural variants of the Yao- and $\Theta$-graphs lead to graphs with $O(fn)$ edges, that are resilient to edge faults of maximum degree $f$ and have stretch factor at most $1+\varepsilon$, for any given $\varepsilon>0$.
1 Introduction
Let $H$ be an undirected “host graph”, in which each edge has a weight. For a real number $t\geq 1$, a subgraph $G$ of $H$ is called a $t$-spanner for $H$, if for any two vertices $p$ and $q$ of $H$,
$\delta_{G}(p,q)\leq t\cdot\delta_{H}(p,q),$
where, for any edge-weighted graph $X$, $\delta_{X}(p,q)$ denotes the length of a shortest path between $p$ and $q$ in $X$. The smallest such value for $t$ is called the stretch factor of $G$.
There is a large amount of literature on the problem of designing algorithms that compute a $t$-spanner for a graph $H$, where the goal is to obtain the best trade-offs between the stretch factor $t$ and the number of edges in $G$. For general edge-weighted graphs $H$, we refer the reader to the survey paper Ahmed et al. [1].
In the metric case, the host graph $H$ is a complete graph $K_{S}$ induced by a finite metric space $(S,|\cdot|)$: The vertex set of $K_{S}$ is $S$ and for any two distinct points $p$ and $q$ in $S$, $K_{S}$ contains the edge $\{p,q\}$ with weight $|pq|$. In this case, a graph $G$ with vertex set $S$ is a $t$-spanner for $S$, if for any two points $p$ and $q$ of $S$,
$\delta_{G}(p,q)\leq t\cdot|pq|.$
For several classes of metric spaces, constructions are known that compute, for any constant $t>1$, a $t$-spanner with $O(n)$ edges, where $n$ is the number of points in $S$. For example, this is the case for Euclidean spaces, where $S$ is a finite set of points in $\mathbb{R}^{d}$ (where the dimension $d$ is a constant) and the distance between two points is their Euclidean distance. This was first shown by Keil [12]. Many other constructions can be found in Narasimhan and Smid [17]. Another example is the case when the metric space $(S,|\cdot|)$ has bounded doubling dimension. For this case, Har-Peled and Mendel [10] were the first to show how to compute a $t$-spanner with $O(n)$ edges. Other results can be found in Filtser and Solomon [8].
1.1 Fault-Tolerant Spanners
Spanners were generalized by Levcopoulos et al. [13, 14] to fault-tolerant spanners. Let $f\geq 1$ be an integer. A subgraph $G$ of the host graph $H$ is an $f$-fault-tolerant $t$-spanner for $H$, if the following holds for any subset $F$ of at most $f$ edges of $G$: For any two vertices $p$ and $q$ of $H$,
$\delta_{G\setminus F}(p,q)\leq t\cdot\delta_{H\setminus F}(p,q),$ (1)
where, for any edge-weighted graph $X=(V_{X},E_{X})$, $X\setminus F$ denotes the graph with vertex set $V_{X}$ and edge set $E_{X}\setminus F$. Thus, $\delta_{G\setminus F}(p,q)$ denotes the length of a shortest path between $p$ and $q$ in the graph obtained by removing all edges in $F$ from $G$. We remark that there are also versions of fault-tolerant spanners in which up to $f$ vertices and edges can be removed.
Again, there is a large amount of literature on the problem of constructing fault-tolerant spanners. For the case of general edge-weighted graphs, the first results were given by Chechik et al. [6]. See Bodwin et al. [2] for more recent results. For Euclidean spaces and metric spaces of bounded doubling dimension, we refer the reader to [5, 11, 15, 16, 17, 18].
Recently, Bodwin, Haeupler, and Parter [3] introduced a more general notion of fault-tolerant spanners. A subgraph $G=(V_{H},E)$ of the host graph $H=(V_{H},E_{H})$ is an $f$-faulty-degree $t$-spanner for $H$, if for any subset $F$ of $E$, for which the maximum degree of the graph $(V_{H},F)$ is at most $f$, and for any two vertices $p$ and $q$ in $V_{H}$, (1) holds.
For example, in a $1$-faulty-degree $t$-spanner, we are allowed to remove a perfect matching that has $|V_{H}|/2$ edges, whereas in a $1$-fault-tolerant $t$-spanner, we are only allowed to remove one edge. This shows that an $f$-faulty-degree spanner is a much stronger condition.
Bodwin, Haeupler, and Parter [3] proved the following result. Let $H$ be any edge-weighted graph with $n$ vertices. For any positive integers $f$ and $k$, there exists an $f$-faulty-degree $(2k-1)$-spanner for $H$ that has $O(k^{k}\cdot f^{1-1/k}\cdot n^{1+1/k})$ edges. Notice that there is a large dependance on $k$. In particular, the upper bound on the number of edges is subquadratic in $n$ only if $k=o(\log n/\log\log n)$. [^1] Recall that $k^{k}=n$ solves to $k\sim\log n/\log\log n$.
1.2 Our Results
We consider faulty-degree spanners for the case when the host graph is the complete graph $K_{S}$ induced by a finite metric space $(S,|\cdot|)$. For completeness, we state the definition of a faulty-degree spanner for this case.
Definition 1
Let $(S,|\cdot|)$ be a finite metric space, let $f\geq 0$ be an integer, let $t\geq 1$ be a real number, and let $G=(S,E)$ be an undirected graph, in which each edge $\{p,q\}$ has weight $|pq|$. The graph $G$ is called an $f$-faulty-degree $t$-spanner for $S$, if the following holds for any subset $F$ of $E$, for which the maximum degree of the graph $(S,F)$ is at most $f$: For any two points $p$ and $q$ in $S$,
$\delta_{G\setminus F}(p,q)\leq t\cdot\delta_{K_{s}\setminus F}(p,q).$
In Section 2, we consider arbitrary metric spaces $(S,|\cdot|)$. We will show the following: Assume that $G^{\prime}=(S,E^{\prime})$ is an arbitrary $t$-spanner for $S$. Let $n$ denote the number of points in $S$, and let $m$ denote the number of edges in $E^{\prime}$. For any integer $f\geq 1$, we can convert $G^{\prime}$ to a graph $G=(S,E)$ with at most $(4f-1)m$ edges, such that $G$ is an $f$-faulty-degree $t^{\prime}$-spanner for $S$, where $t^{\prime}$ is a function of $f$ and $t$. The construction is valid for any $f$ with $21\leq f\leq(n-1)/$ 2. In Section 2.1, we consider the case when $f=1$ and prove that $t^{\prime}\leq 3t$. We also give an example for which $t^{\prime}$ is arbitrarily close to $3t$. In Section 2.2, we consider the general case $f\geq 1$ and prove that $t^{\prime}\leq(8f+2)t$. For $f\geq 3$, we give an example for which $t^{\prime}=\Omega(ft)$.
In Section 3, we assume that $S$ is a set of $n$ points in a metric space that admits a well-separated pair decomposition (WSPD). Let $m$ denote the size of such a WSPD. We will show that, for any integer $f\geq 1$ and any real constant $\varepsilon>0$, there exists an $f$-faulty-degree $(1+\varepsilon)$-spanner for $S$ that has at most $(2f+1)^{2}m$ edges. We remark that our spanner is obtained by modifying the construction in Levcopoulos et al. [14] for standard fault-tolerant spanners. Examples of metric spaces for which $m=O(n)$ are Euclidean spaces of constant dimension (see Callahan and Kosaraju [4]) and metric spaces of bounded doubling dimension (see Har-Peled and Mendel [10]).
In Section 4, we will show that in Euclidean space, we can do even better: Assume that $S$ is a set of $n$ points in $\mathbb{R}^{d}$, where the dimension $d$ is a constant, and the distance $|pq|$ between two points $p$ and $q$ is their Euclidean distance. For this case, we will define two graphs with vertex set $S$ that are variants of the Yao-graph (see Flinchbaugh and Jones [9] and Yao [19]) and the $\Theta$-graph (see Clarkson [7] and Keil [12]) and that depend on the maximum degree bound $f$ of any set $F$ of faulty edges. We will prove that, for any real constant $\varepsilon>0$, each of these graphs is an $f$-faulty-degree $(1+\varepsilon)$-spanner for $S$ with $O(fn)$ edges. In particular, the stretch factor does not depend on $f$. Observe that this is optimal, up to constant factors, because every vertex in an $f$-faulty-degree spanner must have degree at least $f$. We remark that our graphs are variations of the standard fault-tolerant spanners obtained by Lukovszki [15].
We conclude this section by introducing notation that will be used throughout the paper. Let $P=(p_{1},p_{2},\ldots,p_{k})$ be a sequence of points in a finite metric space $(S,|\cdot|)$. The length of this sequence is denoted by $|P|$ and is equal to $\sum_{i=1}^{k-1}|p_{i}p_{i+1}|$. If all points are pairwise distinct, then $P$ is a path in $K_{S}$. If the sequence may have repeated points, then $P$ is a walk in $K_{S}$.
2 General Metric Spaces
In this section, $(S,|\cdot|)$ denotes an arbitrary finite metric space. Let $n$ be the number of points in $S$ and let $t\geq 1$ be a real number. We assume that we are given a $t$-spanner $G^{\prime}=(S,E^{\prime})$ for $S$ and we denote the number of edges of $G^{\prime}$ by $m$. For any integer $f$ with $21\leq f\leq(n-1)/$ 2, we will show how to convert $G^{\prime}$ to a graph $G=(S,E)$ that has at most $(4f-1)m$ edges and is an $f$-faulty-degree $t^{\prime}$-spanner for $S$, where $t^{\prime}$ is a function of $f$ and $t$.
The graph $G=(S,E)$ is obtained as follows (refer to Figure 1):
Figure 1: Illustrating the edges that are added to $E$ because of the edge $\{a,b\}$ in $E^{\prime}$. In this figure, $f=3$.
  • The edge set $E$ contains all edges of $E^{\prime}$.
  • For each edge $\{a,b\}$ in $E^{\prime}$, let $c_{1},c_{2},\ldots,c_{n-2}$ be the sequence of points of $S\setminus\{a,b\}$, in non-decreasing order of the values $|ac_{i}|+|c_{i}b|$; ties are broken arbitrarily. Let $C_{ab}$ denote the set consisting of the first $12f-$ 1 points in this sequence, i.e., $C_{ab}=\{c_{1},c_{2},\ldots,c_{2f-1}\}$. The edge set $E$ contains the $24f-$ 2 edges $\{a,c_{i}\}$ and $\{c_{i},b\}$, for $i=1,2,\ldots,2f-1$.
Observe that the number of edges in $E$ is at most $m+(4f-2)m=(4f-1)m$.
In the rest of this section, we will prove that $G$ is an $f$-faulty-degree $t^{\prime}$-spanner for $S$. In Section 2.1, we consider the case when $f=1$, in which case $t^{\prime}\leq 3t$. In Section 2.2, we consider the general case when $f\geq 1$ and prove that $t^{\prime}\leq(8f+2)t$.
2.1 Faulty Matchings: $f=1$
Consider the graph $G=(S,E)$ defined above for the case when $f=1$. Any subset $F$ of $E$ such that the maximum degree of the graph $(S,F)$ is at most $1$ is a matching in $G$. In other words, $G\setminus F$ is obtained by removing this matching from $G$. We will prove the following result.
Theorem 1
Let $(S,|\cdot|)$ be a finite metric space, let $t\geq 1$ be a real number, and let $G^{\prime}=(S,E^{\prime})$ be a $t$-spanner for $S$ with $m$ edges. Then the graph $G=(S,E)$ is a $1$-faulty-degree $(3t)$-spanner for $S$ that has at most $3m$ edges.
We have already seen that the number of edges in $E$ is at most $3m$. It remains to prove that $G$ is a $1$-faulty-degree $(3t)$-spanner for $S$.
Let $F$ be an arbitrary matching in $G$, and let $p$ and $q$ be two points in $S$. We have to show that
$\delta_{G\setminus F}(p,q)\leq 3t\cdot\delta_{K_{s}\setminus F}(p,q).$
Let $P^{\prime}$ be a shortest path between $p$ and $q$ in the graph $G^{\prime}$. Since $G^{\prime}$ is a $t$-spanner for $S$, the length $|P^{\prime}|$ of $P^{\prime}$ is at most $t|pq|$.
Let $\{a_{1},b_{1}\},\{a_{2},b_{2}\},\ldots,\{a_{k},b_{k}\}$ be the edges on $P^{\prime}$ that are in $F$. We may assume that, if we follow $P^{\prime}$ from $p$ to $q$, the endpoints of these edges are visited in the order $a_{1},b_{1},a_{2},b_{2},\ldots,a_{k},b_{k}$; refer to Figure 2. Since $F$ is a matching, $b_{1}\neq a_{2},b_{2}\neq a_{3},\ldots,b_{k-1}\neq a_{k}$.
Figure 2: $P^{\prime}$ is a shortest path between $p$ and $q$ in the graph $G^{\prime}$. The red edges $\{a_{i},b_{i}\}$, $1\leq i\leq k$, are in the matching $F$. $P$ (in green) is a walk between $p$ and $q$ in the graph $G\setminus F$. Note that any vertex $c_{i}$ may be on the path $P^{\prime}$.
For each $i$ with $1\leq i\leq k$, let $c_{i}$ be the point in $S\setminus\{a_{i},b_{i}\}$ that minimizes $|a_{i}c_{i}|+|c_{i}b_{i}|$ and that was chosen when defining the graph $G$. Since $F$ is a matching, all edges $\{a_{i},c_{i}\}$ and $\{c_{i},b_{i}\}$, for $1\leq i\leq k$, are in $G\setminus F$.
We define $P$ to be the walk (i.e., possibly non-simple path) that
  • starts at $p$ and follows $P^{\prime}$ to $a_{1}$,
  • for $i=1,2,\ldots,k-1$, takes the two edges $\{a_{i},c_{i}\}$ and $\{c_{i},b_{i}\}$, and then follows $P^{\prime}$ from $b_{i}$ to $a_{i+1}$,
  • takes the two edges $\{a_{k},c_{k}\}$ and $\{c_{k},b_{k}\}$, and then follows $P^{\prime}$ from $b_{k}$ to $q$.
Refer to Figure 2. Observe that $P$ is a walk between $p$ and $q$ in the graph $G\setminus F$.
If $k=0$, then $P=P^{\prime}$ and we have
$\delta_{G\setminus F}(p,q)=|P^{\prime}|\leq t|pq|=t\cdot\delta_{K_{S}\setminus F}(p,q)\leq 3t\cdot\delta_{K_{S}\setminus F}(p,q).$
Assume that $k=1$ and $P^{\prime}$ consists of just the edge $\{p,q\}$. Then $p=a_{1}$, $q=b_{1}$, and the length of $P$ is equal to $|a_{1}c_{1}|+|c_{1}b_{1}|$. By the definition of $c_{1}$, we have $\delta_{K_{S}\setminus F}(p,q)=|a_{1}c_{1}|+|c_{1}b_{1}|$ and, thus,
$\delta_{G\setminus F}(p,q)=\delta_{K_{S}\setminus F}(p,q)\leq 3t\cdot\delta_{K_{S}\setminus F}(p,q).$
It remains to consider the case covered by the following lemma.
Lemma 1
Assume that $k\geq 1$ and $P^{\prime}$ has at least two edges. Then the graph $G\setminus F$ contains a walk $Q$ between $p$ and $q$ whose length is at most $3|P^{\prime}|$.
This lemma will imply that
$\delta_{G\setminus F}(p,q)\leq|Q|\leq 3|P^{\prime}|\leq 3t|pq|\leq 3t\cdot\delta_{K_{s}\setminus F}(p,q).$
For any two vertices $x$ and $y$ of $P^{\prime}$, we denote by $P^{\prime}_{xy}$ the subpath of $P^{\prime}$ between $x$ and $y$. Using this notation, the length $|P|$ of $P$ satisfies
$|P|=|P^{\prime}_{pa_{1}}|+\sum_{i=1}^{k-1}|P^{\prime}_{b_{i}a_{i+1}}|+|P^{\prime}_{b_{k}q}|+\sum_{i=1}^{k}\left(|a_{i}c_{i}|+|c_{i}b_{i}|\right).$ (2)
To prove Lemma 1, we consider five cases.
Case 1: The edge of $P^{\prime}$ that is incident on $p$ is not in $F$.
Since $p\not\in\{a_{1},b_{1}\}$, our choice of $c_{1}$ implies that
$|a_{1}c_{1}|+|c_{1}b_{1}|\leq|a_{1}p|+|pb_{1}|.$
For each $i$ with $2\leq i\leq k$, since $b_{i-1}\not\in\{a_{i},b_{i}\}$, our choice of $c_{i}$ implies that
$|a_{i}c_{i}|+|c_{i}b_{i}|\leq|a_{i}b_{i-1}|+|b_{i-1}b_{i}|.$
It follows that
$\sum_{i=1}^{k}\left(|a_{i}c_{i}|+|c_{i}b_{i}|\right)$$\leq$ $\left(|a_{1}p|+|pb_{1}|\right)+\sum_{i=2}^{k}\left(|a_{i}b_{i-1}|+|b_{i-1}b_{i}|\right)$
$\leq$ $|P^{\prime}_{pa_{1}}|+|P^{\prime}_{pb_{1}}|+\sum_{i=2}^{k}\left(|P^{\prime}_{b_{i-1}a_{i}}|+|P^{\prime}_{b_{i-1}b_{i}}|\right)$
$\leq$ $2|P^{\prime}|.$
Using this, it follows from (2) that $|P|\leq 3|P^{\prime}|$. Hence, Lemma 1 holds for $Q=P$.
Case 2: The edge of $P^{\prime}$ that is incident on $q$ is not in $F$.
By switching the roles of $p$ and $q$, this case is the same as Case 1.
Case 3: $k=2$, the path $P^{\prime}$ consists of three edges, the edge of $P^{\prime}$ that is incident on $p$ is in $F$, and the edge of $P^{\prime}$ that is incident on $q$ is in $F$.
In this case, $p=a_{1}$, $q=b_{2}$, and the path $P$ consists of the five edges $\{a_{1},c_{1}\}$, $\{c_{1},b_{1}\}$, $\{b_{1},a_{2}\}$, $\{a_{2},c_{2}\}$, and $\{c_{2},b_{2}\}$. Observe that $\{b_{1},a_{2}\}$ is not in $F$.
Let $c$ be the point in $S\setminus\{b_{1},a_{2}\}$ that minimizes $|b_{1}c|+|ca_{2}|$ and that was chosen when defining the graph $G$.
Case 3.1: $c\not\in\{a_{1},b_{2}\}$.
Since $a_{1}\not\in\{b_{1},a_{2}\}$, our choice of $c$ implies that
$|b_{1}c|+|ca_{2}|\leq|b_{1}a_{1}|+|a_{1}a_{2}|\leq|b_{1}a_{1}|+|a_{1}b_{1}|+|b_{1}a_{2}|.$
Similarly, since $b_{2}\not\in\{b_{1},a_{2}\}$,
$|b_{1}c|+|ca_{2}|\leq|b_{1}b_{2}|+|b_{2}a_{2}|\leq|b_{1}a_{2}|+|a_{2}b_{2}|+|b_{2}a_{2}|.$
By adding these two inequalities, we get
$2\left(|b_{1}c|+|ca_{2}|\right)\leq 2|a_{1}b_{1}|+2|b_{1}a_{2}|+2|a_{2}b_{2}|=2|P^{\prime}|,$
which gives
$|b_{1}c|+|ca_{2}|\leq|P^{\prime}|.$ (3)
Since $c\not\in\{a_{1},b_{1}\}$, our choice of $c_{1}$ implies that
$|a_{1}c_{1}|+|c_{1}b_{1}|\leq|a_{1}c|+|cb_{1}|.$ (4)
Since $c\not\in\{a_{2},b_{2}\}$, our choice of $c_{2}$ implies that
$|a_{2}c_{2}|+|c_{2}b_{2}|\leq|a_{2}c|+|cb_{2}|.$ (5)
Using (4) and (5), we get
$|P|$ $=$ $|a_{1}c_{1}|+|c_{1}b_{1}|+|b_{1}a_{2}|+|a_{2}c_{2}|+|c_{2}b_{2}|$
$\leq$ $|a_{1}c|+|cb_{1}|+|b_{1}a_{2}|+|a_{2}c|+|cb_{2}|.$
Using (3), we get
$|P|$$\leq$ $|a_{1}c|+|b_{1}a_{2}|+|cb_{2}|+|P^{\prime}|$
$\leq$ $\left(|a_{1}b_{1}|+|b_{1}c|\right)+|b_{1}a_{2}|+\left(|ca_{2}|+|a_{2}b_{2}|\right)+|P^{\prime}|$
$=$ $|b_{1}c|+|ca_{2}|+2|P^{\prime}|.$
One more application of (3) gives
$|P|\leq 3|P^{\prime}|.$
Hence, Lemma 1 holds for $Q=P$.
Case 3.2: $c=a_{1}$.
By the definition of the graph $G$, $\{c,a_{2}\}=\{p,a_{2}\}$ is an edge in $G$. Since $F$ is a matching, this edge is in the graph $G\setminus F$. Define the path $Q=(p,a_{2},c_{2},b_{2}=q)$. This path is in $G\setminus F$.
Since $c\not\in\{a_{2},b_{2}\}$, our choice of $c_{2}$ implies that
$|a_{2}c_{2}|+|c_{2}b_{2}|\leq|a_{2}c|+|cb_{2}|,$
i.e.,
$|a_{2}c_{2}|+|c_{2}q|\leq|a_{1}a_{2}|+|a_{1}b_{2}|.$
The length of the path $Q$ satisfies
$|Q|=|pa_{2}|+|a_{2}c_{2}|+|c_{2}q|\leq|pa_{2}|+|a_{1}a_{2}|+|a_{1}b_{2}|.$
By the triangle inequality, we have $|Q|\leq 3|P^{\prime}|$. Hence, Lemma 1 holds for $Q$.
Case 3.3: $c=b_{2}$.
By switching the roles of $p$ and $q$, this case is the same as Case 3.2.
Case 4: $k=2$, the path $P^{\prime}$ consists of at least four edges, the edge of $P^{\prime}$ that is incident on $p$ is in $F$, and the edge of $P^{\prime}$ that is incident on $q$ is in $F$.
In this case, $p=a_{1}$ and $q=b_{2}$. Let $z$ be the third point on $P^{\prime}$, when traversing this path from $p$ to $q$. Then, the edge $\{b_{1},z\}$ is on $P^{\prime}$ and $z\neq a_{1}$.
Observe that $P^{\prime}_{pz}$ is a shortest path between $p$ and $z$ in $G^{\prime}$ and this path satisfies the condition of Case 2. Therefore, the graph $G\setminus F$ contains a walk $Q_{1}$ between $p$ and $z$ whose length is at most $3|P^{\prime}_{pz}|$.
Similarly, $P^{\prime}_{zq}$ is a shortest path between $z$ and $q$ in $G^{\prime}$ and this path satisfies the condition of Case 1. Thus, the graph $G\setminus F$ contains a walk $Q_{2}$ between $z$ and $q$ whose length is at most $|P^{\prime}_{zq}|$.
Let $Q$ be the concatenation of $Q_{1}$ and $Q_{2}$. Then $Q$ is a walk between $p$ and $q$ in $G\setminus F$. The length of this walk satisfies
$|Q|=|Q_{1}|+|Q_{2}|\leq 3|P^{\prime}_{pz}|+3|P^{\prime}_{zq}|=3|P^{\prime}|.$
Thus, Lemma 1 holds for this walk $Q$.
Case 5: $k\geq 3$, the edge of $P^{\prime}$ that is incident on $p$ is in $F$, and the edge of $P^{\prime}$ that is incident on $q$ is in $F$.
In this case, $p=a_{1}$ and $q=b_{k}$. Observe that $P^{\prime}_{pb_{2}}$ is a shortest path between $p$ and $b_{2}$ in $G^{\prime}$ and this path satisfies one of the conditions of Cases 3 and 4. Therefore, the graph $G\setminus F$ contains a walk $Q_{1}$ between $p$ and $b_{2}$ whose length is at most $3|P^{\prime}_{pb_{2}}|$. Similarly, $P^{\prime}_{b_{2}q}$ is a shortest path between $b_{2}$ and $q$ in $G^{\prime}$ and this path satisfies the condition of Case 1. Thus, the graph $G\setminus F$ contains a walk between $b_{2}$ and $q$ whose length is at most $3|P^{\prime}_{b_{2}q}|$. Let $Q$ be the concatenation of $Q_{1}$ and $Q_{2}$. Then $Q$ is a walk between $p$ and $q$ in $G\setminus F$. The length of this walk satisfies
$|Q|=|Q_{1}|+|Q_{2}|\leq 3|P^{\prime}_{pb_{2}}|+3|P^{\prime}_{b_{2}q}|=3|P^{\prime}|.$
Thus, Lemma 1 holds for this walk $Q$.
This completes the proof of Lemma 1 and, therefore, of Theorem 1.
We complete this section by showing that the upper bound of $3t$ on the stretch factor of $G$ is tight. Let $\varepsilon$ be a real number with $0<\varepsilon\ll 1$, and define $\varepsilon^{\prime}=\varepsilon/(4-2\varepsilon)$. Consider the one-dimensional point set $S=\{p,a,b,q\}$, as indicated below. The graph $G^{\prime}$ connects the points in the left-to-right order and is a $1$-spanner for $S$. The graph $G$ contains $G^{\prime}$ and the two edges $\{p,b\}$ and $\{a,q\}$.
Consider the matching $F=\{\{p,a\},\{b,q\}\}$, as indicated in red. We have $\delta_{G\setminus F}(p,q)=3+2\varepsilon^{\prime}$ and $\delta_{K_{S}\setminus F}(p,q)=1+2\varepsilon^{\prime}$. Thus,
$\frac{\delta_{G\setminus F}(p,q)}{\delta_{K_{S}\setminus F}(p,q)}=\frac{3+2\varepsilon^{\prime}}{1+2\varepsilon^{\prime}}=3-\varepsilon,$
i.e., $G^{\prime}$ is a $1$-faulty-degree $(3-\varepsilon)$-spanner for $S$.
2.2 The General Case: $f\geq 1$
Consider the graph $G=(S,E)$ defined in the beginning of Section 2 for the general case when $f\geq 1$. We will prove the following result.
Theorem 2
Let $(S,|\cdot|)$ be a finite metric space, let $n$ be the number of points in $S$, let $t\geq 1$ be a real number, and let $G^{\prime}=(S,E^{\prime})$ be a $t$-spanner for $S$ with $m$ edges. Then, for any integer $f$ with $21\leq f\leq(n-1)/$ 2, the graph $G=(S,E)$ is an $f$-faulty-degree $((8f+2)t)$-spanner for $S$ that has at most $(4f-1)m$ edges.
We have already seen that the number of edges in $E$ is at most $(4f-1)m$. It remains to prove that $G$ is an $f$-faulty-degree $((8f+2)t)$-spanner for $S$.
Let $F$ be an arbitrary subset of the edge set $E$ of $G$, such that the maximum degree of the graph $(S,F)$ is at most $f$.
Lemma 2
The graph $K_{S}\setminus F$ is connected.
Proof. The lemma will follow the following claim: For every partition of $S$ into two non-empty sets $A$ and $B$, the graph $K_{S}\setminus F$ contains an edge between $A$ and $B$.
To prove this claim, we may assume without loss of generality that the size of $B$ is at least the size of $A$. Thus, the size of $B$ is at least $n/2$, which is larger than $f$. Let $p$ be a point in $A$. In $K_{S}$, there are more than $f$ edges of the form $\{p,q\}$, with $q\in B$. At least one these edges is in $K_{S}\setminus F$.
Lemma 3
Let $\{a,b\}$ be an edge in $G^{\prime}$. If this edge is in $F$, then there exists a point $c$ in $C_{ab}$ such that both $\{a,c\}$ and $\{c,b\}$ are edges in the graph $G\setminus F$.
Proof. Assume that for each of the $12f-$ 1 points $c$ in $C_{ab}$, at least one of the two edges $\{a,c\}$ and $\{c,b\}$ is in $F$. Then, in $(S,F)$, the sum of the degrees of $a$ and $b$ is at least $12+(2f-1)=2f+$ 1. Therefore, $a$ or $b$ has degree more than $f$ in $(S,F)$, which is a contradiction.
Lemma 4
Let $\{a,b\}$ be an edge in $G^{\prime}$. Then there exists a shortest path between $a$ and $b$ in the graph $G\setminus F$, all of whose vertices belong to $C_{ab}\cup\{a,b\}$.
Proof. First assume that $\{a,b\}$ is not an edge in $F$. Then this edge is in the graph $G\setminus F$. Using the triangle inequality, this edge is a shortest path between $a$ and $b$. Thus, the lemma holds.
Figure 3: Illustrating the proof of Lemma 4.
Now assume that $\{a,b\}$ is in $F$. Let $c$ be a point in $C_{ab}$ as in Lemma 3. The path $P=(a,c,b)$ is in $G\setminus F$. Let $Q$ be a shortest path between $a$ and $b$ in $G\setminus F$, and assume that $Q$ contains a vertex $v$ that is not in $C_{ab}\cup\{a,b\}$; refer to Figure 3. Then
$|P|=|ac|+|cb|\leq|av|+|vb|\leq|Q|$
and, thus, $|P|=|Q|$. The path $P$ satisfies the lemma.
Lemma 5
Let $p$ and $q$ be two distinct points in $S$, and let $P^{\prime}$ be a shortest path between $p$ and $q$ in the graph $G^{\prime}$. If the number of vertices of $P^{\prime}$ is at least $22f+$ 2, then
$\delta_{G\setminus F}(p,q)\leq(8f+2)\cdot|P^{\prime}|.$
Proof. Let $e=\{a,b\}$ be any edge on the path $P^{\prime}$. By Lemma 4, there exists a shortest path $P_{e}$ between $a$ and $b$ in the graph $G\setminus F$, all of whose vertices belong to $C_{ab}\cup\{a,b\}$; refer to Figure 4.
Figure 4: The black path $P^{\prime}$ is a shortest path between $p$ and $q$ in $G^{\prime}$. For each edge $e=\{a,b\}$ on $P^{\prime}$, the green path $P_{e}$ is a shortest path between $a$ and $b$ in $G\setminus F$. The concatenation of all paths $P_{e}$ is a walk $P$ between $p$ and $q$ in $G\setminus F$.
Let $P$ be the concatenation of all paths $P_{e}$, where $e$ ranges over all edges of $P^{\prime}$. Then $P$ is a walk between $p$ and $q$ in $G\setminus F$. We will prove that the length of $P$ is at most $(8f+2)\cdot|P^{\prime}|$. This will prove the lemma.
Figure 5: $v_{e}$ is a vertex of $P^{\prime}$ that is closest (along $P^{\prime}$ and in terms of the number of edges) to $a$ or $b$, and that is not in $C_{ab}\cup\{a,b\}$.
Again, let $e=\{a,b\}$ be any edge on the path $P^{\prime}$. Recall that the set $C_{ab}\cup\{a,b\}$ has size $12f+$ 1 and $P^{\prime}$ contains at least $22f+$ 2 vertices. Therefore, there exists a vertex of $P^{\prime}$ that is not in $C_{ab}\cup\{a,b\}$. Let $v_{e}$ be such a vertex that is closest (along $P^{\prime}$ and in terms of the number of edges), to $a$ or $b$; refer to Figure 5. Then, again using the fact that the size of $C_{ab}\cup\{a,b\}$ is $12f+$ 1,
  • the subpath $P^{\prime}_{av_{e}}$ of $P^{\prime}$ between $a$ and $v_{e}$ has at most $12f+$ 1 edges, and
  • the subpath $P^{\prime}_{v_{e}b}$ of $P^{\prime}$ between $b$ and $v_{e}$ has at most $12f+$ 1 edges.
First assume that $e=\{a,b\}$ is an edge in $F$. Let $c$ be a point as in Lemma 3. Since both edges $\{a,c\}$ and $\{c,b\}$ are in $G\setminus F$, we have
$\delta_{G\setminus F}(a,b)\leq|ac|+|cb|.$
Since $v_{e}\not\in C_{ab}$, we have
$|ac|+|cb|\leq|av_{e}|+|v_{e}b|.$
By the triangle inequality, we have
$|av_{e}|+|v_{e}b|\leq|P^{\prime}_{av_{e}}|+|P^{\prime}_{v_{e}b}|.$
By combining the above three inequalities, we get
$\delta_{G\setminus F}(a,b)\leq|P^{\prime}_{av_{e}}|+|P^{\prime}_{v_{e}b}|.$ (6)
Observe that (6) also holds if $e=\{a,b\}$ is not an edge in $F$, because in that case, $\delta_{G\setminus F}(a,b)=|ab|$, which is at most equal to the right-hand side in (6).
Since $P$ is the concatenation of all paths $P_{e}$, with $e$ an edge of $P^{\prime}$, we have
$|P|$ $=$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}|P_{e}|$ (7)
$=$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}\delta_{G\setminus F}(a,b)$
$\leq$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}\left(|P^{\prime}_{av_{e}}|+|P^{\prime}_{v_{e}b}|\right)$
$=$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}|P^{\prime}_{av_{e}}|+\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}|P^{\prime}_{v_{e}b}|.$
Consider the first summation in (7). For each edge $\{x,y\}$ on the path $P^{\prime}$, let $N_{xy}$ be the number of edges $e=\{a,b\}$ on $P^{\prime}$ such that $\{x,y\}$ is an edge on the subpath $P^{\prime}_{av_{e}}$. Then
$\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}|P^{\prime}_{av_{e}}|=\sum_{\{x,y\}\mbox{\scriptsize{ on }}P^{\prime}}N_{xy}\cdot|xy|.$
To obtain an upper bound on $N_{xy}$, we assume, without loss of generality, that $x$ is on the subpath of $P^{\prime}$ between $p$ and $y$. Define the following subpath $Q_{xy}$ of $P^{\prime}$:
  • $\{x,y\}$ is an edge of $Q_{xy}$.
  • Starting at $x$, traverse $P^{\prime}$ towards $p$ and stop as soon as $2f$ edges have been traversed or the point $p$ has been reached. All edges traversed are edges of $Q_{xy}$.
  • Starting at $y$, traverse $P^{\prime}$ towards $q$ and stop as soon as $2f$ edges have been traversed or the point $q$ has been reached. All edges traversed are edges of $Q_{xy}$.
Recall that each subpath $P^{\prime}_{av_{e}}$ has at most $12f+$ 1 edges. It follows that every edge $e=\{a,b\}$ that is counted in $N_{xy}$ is an edge on $Q_{xy}$. Since $Q_{xy}$ has at most $14f+$ 1 edges, it follows that $N_{xy}\leq 4f+1$. Thus,
$\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}|P^{\prime}_{av_{e}}|\leq(4f+1)\sum_{\{x,y\}\mbox{\scriptsize{ on }}P^{\prime}}|xy|=(4f+1)\cdot|P^{\prime}|.$
By a symmetric argument, the second summation in (7) is at most $(4f+1)\cdot|P^{\prime}|$. We conclude that $|P|\leq(8f+2)|P^{\prime}|$.
The next lemma will complete the proof of Theorem 2.
Lemma 6
Let $p$ and $q$ be two distinct points in $S$. Then
$\delta_{G\setminus F}(p,q)\leq(8f+2)t\cdot\delta_{K_{S}\setminus F}(p,q).$
Proof. Let $Q$ be a shortest path between $p$ and $q$ in $K_{S}\setminus F$. By Lemma 2, $Q$ exists. For each edge $e=\{a,b\}$ on $Q$, let $P^{\prime}_{e}$ be a shortest path between $a$ and $b$ in $G^{\prime}$. Since $G^{\prime}$ is a $t$-spanner for $S$, we have $|P^{\prime}_{e}|\leq t|ab|$. We will show that
$\delta_{G\setminus F}(a,b)\leq(8f+2)\cdot|P^{\prime}_{e}|.$ (8)
This will imply that
$\delta_{G\setminus F}(p,q)$$\leq$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}Q}\delta_{G\setminus F}(a,b)$
$\leq$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}Q}(8f+2)\cdot|P^{\prime}_{e}|$
$\leq$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}Q}(8f+2)t\cdot|ab|$
$=$ $(8f+2)t\cdot|Q|$
$=$ $(8f+2)t\cdot\delta_{K_{S}\setminus F}(p,q).$
If $P^{\prime}_{e}$ has at least $22f+$ 2 vertices, then (8) follows from Lemma 5. From now on, we assume that $P^{\prime}_{e}$ has at most $12f+$ 1 vertices and, thus, at most $2f$ edges.
Since $Q$ is a path in $K_{S}\setminus F$ and $e=\{a,b\}$ is an edge on $Q$, this edge is not in $F$. If $e$ is an edge in $G\setminus F$, then
$\delta_{G\setminus F}(a,b)=|ab|\leq|P^{\prime}_{e}|$
and, thus (8) holds.
It remains to consider the case when $e=\{a,b\}$ is not an edge in $G\setminus F$. Consider the edge $\{a,d\}$ on $P^{\prime}_{e}$ that is incident on $a$. Observe that $d\neq b$, because $\{a,b\}$ is not an edge in $G^{\prime}$.
We first show that for every point $z$ in $C_{ad}$,
$|az|+|zd|\leq|ab|+|bd|.$ (9)
The proof is by contradiction: If $|ab|+|bd|<|az|+|zd|$, then $b$ is an element of $C_{ad}$ and, therefore, by the definition of the graph $G$, $\{a,b\}$ is an edge in $G$, which is a contradiction.
Next, we show that
$\delta_{G\setminus F}(a,d)\leq 2\cdot|P^{\prime}_{e}|.$ (10)
If $\{a,d\}$ is not in $F$, then $\delta_{G\setminus F}(a,d)=|ad|\leq|P^{\prime}_{e}|$ and, thus, (10) holds. Assume that $\{a,d\}$ is an edge in $F$. By Lemma 3, there exists a point $c$ in $C_{ad}$, such that both $\{a,c\}$ and $\{c,d\}$ are edges in $G\setminus F$. Then, $\delta_{G\setminus F}(a,d)\leq|ac|+|cd|$, which, by (9), is at most $|ab|+|bd|$, which is at most $2\cdot|P^{\prime}_{e}|$, i.e., (10) holds.
Let $D=C_{ad}\cup\{a,d\}$. Observe that $D$ has size $12f+$ 1. Let $\{x,y\}$ be an arbitrary edge on $P^{\prime}_{e}$ with $\{x,y\}\neq\{a,d\}$. Let $D^{\prime}=D\setminus\{x,y\}$. Then the size of $D^{\prime}$ is at least $12f-$ 1. Let $d^{\prime}$ be a point in $D^{\prime}$ for which $|xd^{\prime}|+|d^{\prime}y|$ is maximum. Refer to Figure 6.
Figure 6: $D=C_{ad}\cup\{a,d\}$, $\{x,y\}$ is an edge on $P^{\prime}_{e}$, $\{x,y\}\neq\{a,d\}$.
First, we assume that $\{x,y\}$ is an edge in $F$. By Lemma 3, there is a point $c^{\prime}$ in $C_{xy}$ such that both $\{x,c^{\prime}\}$ and $\{c^{\prime},y\}$ are edges in $G\setminus F$. It is clear that
$\delta_{G\setminus F}(x,y)\leq|xc^{\prime}|+|c^{\prime}y|.$
We next argue that
$|xc^{\prime}|+|c^{\prime}y|\leq|xd^{\prime}|+|d^{\prime}y|.$
Assume, by contradiction, that $|xd^{\prime}|+|d^{\prime}y|<|xc^{\prime}|+|c^{\prime}y|$. Then for all $z$ in $D^{\prime}=D\setminus\{x,y\}$, $|xz|+|zy|<|xc^{\prime}|+|c^{\prime}y|$. There are at least $12f-$ 1 such points $z$, and each of them is in $S\setminus\{x,y\}$. Thus, $c^{\prime}\not\in C_{xy}$, which is a contradiction.
The previous two inequalities, together with the triangle inequality, imply that
$\delta_{G\setminus F}(x,y)\leq|xd^{\prime}|+|d^{\prime}y|$$\leq$ $\left(|P^{\prime}_{xa}|+|ad^{\prime}|\right)+\left(|d^{\prime}d|+|P^{\prime}_{dy}|\right)$
$\leq$ $|ad^{\prime}|+|d^{\prime}d|+2\cdot|P^{\prime}_{e}|.$
We claim that
$|ad^{\prime}|+|d^{\prime}d|\leq|ab|+|bd|.$ (11)
If $d^{\prime}\in\{a,d\}$, then (11) follows from the triangle inequality. If $d^{\prime}\not\in\{a,d\}$, then $d^{\prime}\in C_{ad}$ and (11) follows from (9).
Since $|ab|+|bd|\leq 2\cdot|P^{\prime}_{e}|$, we conclude that
$\delta_{G\setminus F}(x,y)\leq 4\cdot|P^{\prime}_{e}|.$ (12)
In the derivation of (12), we assumed that $\{x,y\}$ is an edge in $F$. If $\{x,y\}$ is not an edge in $F$, then (12) also holds, because, in that case, $\delta_{G\setminus F}(x,y)=|xy|\leq|P^{\prime}_{e}|$.
For each edge $e^{\prime}=\{x,y\}$ on $P^{\prime}_{e}$, let $P_{e^{\prime}}$ be a shortest path between $x$ and $y$ in $G\setminus F$. Let $P_{ab}$ be the concatenation of all paths $P_{e^{\prime}}$, where $e^{\prime}$ ranges over all edges of $P^{\prime}_{e}$. Observe that $P_{ab}$ is a walk between $a$ and $b$ in $G\setminus F$. We have
$\delta_{G\setminus F}(a,b)$$\leq$ $|P_{ab}|$
$=$ $\sum_{e^{\prime}=\{x,y\}\mbox{\scriptsize{ on }}P^{\prime}_{e}}|P_{e^{\prime}}|$
$=$ $\sum_{e^{\prime}=\{x,y\}\mbox{\scriptsize{ on }}P^{\prime}_{e}}\delta_{G\setminus F}(x,y)$
$\leq$ $\sum_{e^{\prime}=\{x,y\}\mbox{\scriptsize{ on }}P^{\prime}_{e}}4\cdot|P^{\prime}_{e}|.$
Since $P^{\prime}_{e}$ has at most $2f$ edges, the latter summation has at most $2f$ terms. We conclude that
$\delta_{G\setminus F}(a,b)\leq 8f\cdot|P^{\prime}_{e}|$
and, thus, (8) holds.
2.2.1 A Lower Bound for our Construction
We have shown that any $t$-spanner $G^{\prime}$ can be transformed to an $f$-faulty-degree $t^{\prime}$-spanner with $t^{\prime}=O(ft)$. Below, we will give an example for which $t^{\prime}=\Omega(ft)$.
Let $f\geq 3$ be an integer and let $\varepsilon$ be a real number with $0<\varepsilon\ll 1/(f-2)$. Let $B_{0},B_{1},\ldots B_{f}$ be pairwise disjoint sets, where $B_{0}=\{b_{0},b_{1},\ldots,b_{f}\}$, and each $B_{i}$, for $1\leq i\leq f$, consists of $2f$ points. Let $S$ be the union of $B_{0},B_{1},\ldots B_{f}$.
We define the following graph $H$ with vertex set $S$; refer to Figure 7:
Figure 7: Illustrating the graph $H$. All black edges have weight $1$, all blue edges have weight $\varepsilon$, and the dotted edge $\{b_{0},b_{f}\}$ has weight $1+\varepsilon$. The set $B_{0}$ is equal to $\{b_{0},b_{1},\ldots,b_{f}\}$. The graph $G^{\prime}$ consists of all solid edges.
  • $H$ contains the edge $\{b_{0},b_{f}\}$ having weight $1+\varepsilon$.
  • $H$ contains the edges $\{b_{0},b_{1}\}$ and $\{b_{f-1},b_{f}\}$, each having weight $1$.
  • $H$ contains the edges $\{b_{j},b_{j+1}\}$, for $j=1,2,\ldots,f-2$, each having weight $\varepsilon$.
  • For each $i$ with $1\leq i\leq f$, $H$ contains the complete graph on $B_{i}$, with each edge having weight $\varepsilon$.
  • For each $i$ with $1\leq i\leq f$, $H$ contains the complete bipartite graph on $B_{i}\times\{b_{i-1},b_{i}\}$, with each edge having weight $1$.
For any two points $p$ and $q$ of $S$, we define $|pq|$ to be the shortest-path distance between $p$ and $q$ in the graph $H$. It is clear that $(S,|\cdot|)$ is a metric space.
Observation 1
For each edge $\{p,q\}$ in $H$, $|pq|$ is the weight of this edge, i.e., $\{p,q\}$ is the shortest path between $p$ and $q$ in $H$.
Proof. The proof can be obtained by verifying that for each edge $\{p,q\}$, any path between $p$ and $q$ that has at least two edges has length more than the weight of $\{p,q\}$.
Let $G^{\prime}$ be the graph obtained from $H$ by removing the edge $\{b_{0},b_{f}\}$.
Observation 2
$G^{\prime}$ is a $3$-spanner for $S$.
Proof. If $\{x,y\}$ is an edge of $H$ and $\{x,y\}\neq\{b_{0},b_{f}\}$, then, by Observation 1, $\delta_{G^{\prime}}(x,y)=|xy|$. For the edge $\{b_{0},b_{f}\}$, we have, using the fact that $\varepsilon<1/(f-2)$,
$\delta_{G^{\prime}}(b_{0},b_{f})=1+(f-2)\varepsilon+1\leq 3\leq 3\cdot|b_{0}b_{f}|.$
It follows that, for any two points $p$ and $q$ in $S$,
$\delta_{G^{\prime}}(p,q)\leq 3\cdot\delta_{H}(p,q)=3\cdot|pq|.$
Having defined the $3$-spanner $G^{\prime}$, consider the graph $G$ obtained from $G^{\prime}$ by applying the transformation given in the beginning of Section 2.2.
Observation 3
$\{b_{0},b_{f}\}$ is not an edge in the graph $G$.
Proof. Let $\{p,q\}$ be any edge in $G^{\prime}$. Recall that $G$ contains the complete graph with vertex set $C_{pq}\cup\{p,q\}$. The claim holds if we can show that at least one of $b_{0}$ and $b_{f}$ is not in $C_{pq}\cup\{p,q\}$.
Observe that $|pb_{0}|+|b_{0}q|\geq 2+\varepsilon$ or $|pb_{f}|+|b_{f}q|\geq 2+\varepsilon$, while there are at least $2f$ points $x$ in $S\setminus\{p,q\}$ such that $|px|+|xq|\leq 2$. This implies that at least one of $b_{0}$ and $b_{f}$ is not in $C_{pq}\cup\{p,q\}$.
Observation 4
For any $i$ and $j$ with $1\leq i<j\leq f$, the graph $G$ does not contain any edge between the sets $B_{i}$ and $B_{j}$.
Proof. For any point $a$ in $B_{i}$ and any point $b$ in $B_{j}$, we have $|ab|\geq 2$. For any edge $\{p,q\}$ in $G^{\prime}$, we have $|pa|+|aq|\geq 2+\varepsilon$ or $|pb|+|bq|\geq 2+\varepsilon$, while there are at least $2f$ points $x$ in $S\setminus\{p,q\}$ such that $|px|+|xq|\leq 2$. This implies that at least one of $a$ and $b$ is not in $C_{pq}\cup\{p,q\}$.
By Theorem 2, the graph $G$ is an $f$-faulty-degree $t^{\prime}$-spanner for $S$, where $t^{\prime}\leq 3(8f+2)=24f+6$. Below, we show that $t^{\prime}\geq f$.
Let $F$ be the set of all edges $\{p,q\}$ of $G$ for which both $p$ and $q$ are in $B_{0}$. Since $B_{0}$ has size $f+1$, the maximum degree of the graph $(S,F)$ is at most $f$.
By Observation 3, $\{b_{0},b_{f}\}$ is not an edge in $G\setminus F$. This, together with Observation 4, implies that the shortest path between $b_{0}$ and $b_{f}$ in $G\setminus F$ consists of the following:
  • For $i=0,1,\ldots,f-1$, it takes the edge from $b_{i}$ to an arbitrary point, say $x_{i+1}$, of $B_{i+1}$, and then the edge from $x_{i+1}$ to $b_{i+1}$.
Thus,
$\delta_{G\setminus F}(b_{0},b_{f})=2f.$
Since by Observation 3, $\{b_{0},b_{f}\}$ is not in $F$, we have
$\delta_{K_{S}\setminus F}(b_{0},b_{f})=|b_{0}b_{f}|=1+\varepsilon\leq 2.$
It follows that
$\frac{\delta_{G\setminus F}(b_{0},b_{f})}{\delta_{K_{S}\setminus F}(b_{0},b_{f})}\geq f.$
3 Metric Spaces with a Small WSPD
The construction in the previous section is valid for any metric space that admits a $t$-spanner, for some value of $t$. However, the stretch factor of the $f$-faulty-degree spanner is $\Theta(ft)$. In this section, we will show that, if the metric space admits a well-separated pair decomposition, we can obtain a stretch factor of $1+\varepsilon$, for any given $\varepsilon>0$.
3.1 Preliminaries
Let $(S,|\cdot|)$ be a finite metric space. For any two non-empty subsets $A$ and $B$ of $S$, we define their distance $|AB|$ as
$|AB|=\min\{|pq|:p\in A,q\in B\}$
[…]
For a real number $c>0$, called the separation ratio, we say that $A$ and $B$ are well-separated, if
[…]
Lemma 7
Let $c>0$ be a real number, let $A$ and $B$ be two non-empty subsets of $S$ that are well-separated with respect to $c$, let $p$ and $p^{\prime}$ be any two points in $A$, and let $q$ and $q^{\prime}$ be any two points in $B$. Then
  1. 1.
    $|pp^{\prime}|\leq(1/c)|pq|$,
  2. 2.
    $|qq^{\prime}|\leq(1/c)|pq|$, and
  3. 3.
    $|p^{\prime}q^{\prime}|\leq(1+2/c)|pq|$.
Proof. The first claim follows from the following chain of inequalities:
[…]
$\leq$ $(1/c)|AB|$
$\leq$ $(1/c)|pq|.$
The proof of the second claim is symmetric. For the proof of the third claim, we combine the first two claims and use the triangle inequality:
$|p^{\prime}q^{\prime}|$$\leq$ $|p^{\prime}p|+|pq|+|qq^{\prime}|$
$\leq$ $(1/c)|pq|+|pq|+(1/c)|pq|$
$=$ $(1+2/c)|pq|.$
Definition 2 (Callahan and Kosaraju [4])
Let $(S,|\cdot|)$ be a finite metric space and let $c>0$ be a real number. A well-separated pair decomposition (WSPD) for $S$ is a sequence
$\{A_{1},B_{1}\},\{A_{2},B_{2}\},\ldots,\{A_{m},B_{m}\}$
of pairs of non-empty subsets of $S$, for some integer $m$, such that
  1. 1.
    for each $i$ with $1\leq i\leq m$, $A_{i}$ and $B_{i}$ are well-separated, and
  2. 2.
    for any two distinct points $p$ and $q$ of $S$, there is exactly one index $i$ such that
    1. (a)
      $p\in A_{i}$ and $q\in B_{i}$, or
    2. (b)
      $p\in B_{i}$ and $q\in A_{i}$.
The integer $m$ is called the size of the well-separated pair decomposition.
3.2 The Faulty-Degree Spanner
Let $(S,|\cdot|)$ be a finite metric space, let $f\geq 1$ be an integer, and let $\varepsilon>0$ be a real number. Assume that
$\{A_{1},B_{1}\},\{A_{2},B_{2}\},\ldots,\{A_{m},B_{m}\}$
is a well-separated pair decomposition for $S$ with separation ratio
$c=2+4/\varepsilon.$
We define a graph $G=(S,E)$ whose edge set $E$ is obtained as follows: For each $i$ with $1\leq i\leq m$, define the following two subsets $A^{\prime}_{i}$ and $B^{\prime}_{i}$ of $A_{i}$ and $B_{i}$, respectively:
  • If $A_{i}$ has size at most $12f+$ 1, then $A^{\prime}_{i}=A_{i}$.
  • If $A_{i}$ has size at least $22f+$ 2, then $A^{\prime}_{i}$ is an arbitrary subset of size $12f+$ 1 of $A_{i}$.
  • If $B_{i}$ has size at most $12f+$ 1, then $B^{\prime}_{i}=B_{i}$.
  • If $B_{i}$ has size at least $22f+$ 2, then $B^{\prime}_{i}$ is an arbitrary subset of size $12f+$ 1 of $B_{i}$.
The edge set $E$ contains the edges of the complete bipartite graph with vertex sets $A^{\prime}_{i}$ and $B^{\prime}_{i}$.
It is clear that the total number of edges in $E$ is at most $(2f+1)^{2}m$. In the following subsection, we will prove that $G$ is an $f$-faulty-degree $(1+\varepsilon)$-spanner for $S$.
3.3 Analysis
Let $F$ be an arbitrary subset of the edge set of $G$ such that the maximum degree of the graph $(S,F)$ is at most $f$.
Lemma 8
For each edge $\{p,q\}$ in $K_{S}\setminus F$, we have
$\delta_{G\setminus F}(p,q)\leq(1+\varepsilon)\cdot|pq|.$
Proof. Let $i$ be the index in $\{1,2,\ldots,m\}$ such that (a) $p\in A_{i}$ and $q\in B_{i}$ or (b) $p\in B_{i}$ and $q\in A_{i}$. We may assume without loss of generality that (a) holds.
The proof is by induction on the rank of the distance $|pq|$ in the set of all edge lengths in $K_{S}\setminus F$.
For the base case, assume that $\{p,q\}$ is a shortest edge in $K_{S}\setminus F$. We first show that the size of the set $A_{i}$ is at most $2f$. Assume that this set has at least $12f+$ 1 points. Then $K_{S}$ contains at least $12f\geq f+$ 1 edges $\{p,p^{\prime}\}$ for which $p^{\prime}\in A_{i}$. At least one of these edges is in the graph $K_{S}\setminus F$. Since by Lemma 7, $|pp^{\prime}|<|pq|$, this is a contradiction. By a symmetric argument, we can show that the size of the set $B_{i}$ is at most $2f$. It follows that $p\in A_{i}=A^{\prime}_{i}$ and $q\in B_{i}=B^{\prime}_{i}$. Thus, $\{p,q\}$ is an edge in $G$. Since $\{p,q\}\not\in F$, it follows that
$\delta_{G\setminus F}(p,q)=|pq|\leq(1+\varepsilon)\cdot|pq|.$
For the induction step, assume that $\{p,q\}$ is not a shortest edge in $K_{S}\setminus F$. If $\{p,q\}$ is in $G$, then it is in $G\setminus F$ and the lemma holds. Assume from now on that $\{p,q\}$ is not an edge in $G$. Then $p\not\in A^{\prime}_{i}$ or $q\not\in B^{\prime}_{i}$. We distinguish three cases.
Case 1: $p\not\in A^{\prime}_{i}$ and $q\not\in B^{\prime}_{i}$.
We observe that each of $A_{i}$ and $B_{i}$ has at least $22f+$ 2 points, and each of $A^{\prime}_{i}$ and $B^{\prime}_{i}$ has exactly $12f+$ 1 points.
Since $K_{S}$ contains $12f+1\geq f+$ 1 edges $\{p,p^{\prime}\}$ for which $p^{\prime}$ is in $A^{\prime}_{i}$, there is a point $p^{\prime}$ in $A^{\prime}_{i}$ such that $\{p,p^{\prime}\}$ is in $K_{S}\setminus F$. Since, by Lemma 7, $|pp^{\prime}|\leq(1/c)|pq|<|pq|$, we have, by induction,
$\delta_{G\setminus F}(p,p^{\prime})\leq(1+\varepsilon)\cdot|pp^{\prime}|.$
Note that, by the definition of $G$, for each point $q^{\prime}$ in $B^{\prime}_{i}$, the edge $\{p^{\prime},q^{\prime}\}$ is in $G$. Define
$B^{\prime\prime}_{i}=\{q^{\prime}\in B^{\prime}_{i}:\{p^{\prime},q^{\prime}\}\mbox{ is an edge in }G\setminus F\}.$
Since the size of $B^{\prime}_{i}$ is $12f+$ 1, the size of $B^{\prime\prime}_{i}$ is at least $f+1$. Thus, $K_{S}$ contains at least $f+1$ edges $\{q,q^{\prime}\}$ for which $q^{\prime}$ is in $B^{\prime\prime}_{i}$. It follows that there is a point $q^{\prime}$ in $B^{\prime\prime}_{i}$ such that $\{q,q^{\prime}\}$ is in $K_{S}\setminus F$. Since, by Lemma 7, $|qq^{\prime}|\leq(1/c)|pq|<|pq|$, we have, by induction,
$\delta_{G\setminus F}(q,q^{\prime})\leq(1+\varepsilon)\cdot|qq^{\prime}|.$
Note that, since $q^{\prime}$ is in $B^{\prime\prime}_{i}$, $\{p^{\prime},q^{\prime}\}$ is an edge in $G\setminus F$. It follows that
$\delta_{G\setminus F}(p,q)$$\leq$ $\delta_{G\setminus F}(p,p^{\prime})+|p^{\prime}q^{\prime}|+\delta_{G\setminus F}(q^{\prime},q)$
$\leq$ $(1+\varepsilon)\cdot|pp^{\prime}|+|p^{\prime}q^{\prime}|+(1+\varepsilon)\cdot|q^{\prime}q|.$
Using Lemma 7, we conclude that
$\delta_{G\setminus F}(p,q)$$\leq$ $(1+\varepsilon)\cdot(1/c)|pq|+(1+2/c)|pq|+(1+\varepsilon)\cdot(1/c)|pq|$
$=$ $(1+(4+2\varepsilon)/c)|pq|$
$=$ $(1+\varepsilon)|pq|,$
where the last equality follows from our choice for $c$.
Case 2: $p\in A^{\prime}_{i}$ and $q\not\in B^{\prime}_{i}$.
In this case, $B_{i}$ has at least $22f+$ 2 points and $B^{\prime}_{i}$ has exactly $12f+$ 1 points. By the definition of $G$, for each point $q^{\prime}$ in $B^{\prime}_{i}$, the edge $\{p,q^{\prime}\}$ is in $G$. Define
$B^{\prime\prime}_{i}=\{q^{\prime}\in B^{\prime}_{i}:\{p,q^{\prime}\}\mbox{ is an edge in }G\setminus F\}.$
Note that the size of $B^{\prime\prime}_{i}$ is at least $f+1$. Thus, $K_{S}$ contains at least $f+1$ edges $\{q,q^{\prime}\}$ for which $q^{\prime}$ is in $B^{\prime\prime}_{i}$. It follows that there is a point $q^{\prime}$ in $B^{\prime\prime}_{i}$ such that $\{q,q^{\prime}\}$ is in $K_{S}\setminus F$. By Lemma 7, $|qq^{\prime}|<|pq|$, implying that
$\delta_{G\setminus F}(q^{\prime},q)\leq(1+\varepsilon)\cdot|q^{\prime}q|.$
Since $q^{\prime}$ is in $B^{\prime\prime}_{i}$, $\{p,q^{\prime}\}$ is an edge in $G\setminus F$. It follows that
$\delta_{G\setminus F}(p,q)\leq|pq^{\prime}|+\delta_{G\setminus F}(q^{\prime},q)\leq|pq^{\prime}|+(1+\varepsilon)\cdot|q^{\prime}q|.$
As in Case 1, it follows that
$\delta_{G\setminus F}(p,q)\leq(1+\varepsilon)\cdot|pq|.$
Case 3: $p\not\in A^{\prime}_{i}$ and $q\in B^{\prime}_{i}$.
This case is symmetric to Case 2.
We summarize the result of this section in the following theorem:
Theorem 3
Let $(S,|\cdot|)$ be a finite metric space, let $f\geq 1$ be an integer, and let $\varepsilon>0$ be a real number. Assume there exists a well-separated pair decomposition for $S$ with separation ratio $c=2+4/\varepsilon$. Let $m$ be the size of this decomposition. Then the graph $G=(S,E)$ defined at the beginning of Section 3.2 is an $f$-faulty-degree $(1+\varepsilon)$-spanner for $S$ that has at most $(2f+1)^{2}m$ edges.
Proof. Let $F$ be an arbitrary subset of the edge set of $G$ such that the maximum degree of the graph $(S,F)$ is at most $f$. Let $p$ and $q$ be two distinct points in $S$. To prove that
$\delta_{G\setminus F}(p,q)\leq(1+\varepsilon)\cdot\delta_{K_{S}\setminus F}(p,q),$
we consider a shortest path in $K_{S}\setminus F$ and apply Lemma 8 to each edge on this path.
To give two examples, Callahan and Kosaraju [4] have shown that for any set $S$ of $n$ points in $\mathbb{R}^{d}$, where the dimension $d$ is a constant and ||$\cdot$|| ⋅ | is the Euclidean distance function, there is a WSPD for $S$ of size $m=O(n)$. Har-Peled and Mendel [10] have shown that for every metric space of bounded doubling dimension, there is a WSPD for $S$ of size $m=O(n)$. Theorem 3 implies that in both of these examples, we can construct an $f$-faulty-degree $(1+\varepsilon)$-spanner with $O(f^{2}n)$ edges. In the next section, we will improve the number of edges to $O(fn)$ for the Euclidean case.
4 Euclidean Metric Spaces of Constant Dimension
In this section, we consider the metric space $(S,|\cdot|)$, where $S$ is a set of $n$ points in $\mathbb{R}^{d}$, for some constant $d$, and ||$\cdot$|| ⋅ | is the Euclidean distance function.
We will define two graphs with vertex set $S$ that are variants of the Yao-graph (see [9, 19]) and the $\Theta$-graph (see [7, 12]) and that depend on the maximum degree bound $f$ of any set $F$ of faulty edges. As we will see, for any real constant $\varepsilon>0$, each of these graphs is an $f$-faulty-degree $(1+\varepsilon)$-spanner with $O(fn)$ edges, where $n$ is the size of $S$.
4.1 Cones
Let $\theta$ be a real number with $40<\theta<\pi/$ 4 and let $\mathcal{C}$ be a collection of cones that cover $\mathbb{R}^{d}$, such that for each cone $C$ in $\mathcal{C}$, the apex of $C$ is at the origin 00 and the angular diameter of $C$ is at most $\theta$, i.e.,
$\max\left\{\angle(0x,0y):x,y\in C\setminus\{0\}\right\}\leq\theta.$
Lukovszki [16] has shown how to obtain such a collection $\mathcal{C}$ consisting of $O(1/\theta^{d-1})$ cones; see also [17, Chapter 5].
4.2 Yao- and Theta-Graphs
Let $S$ be a set of $n$ points in $\mathbb{R}^{d}$ and let $f\geq 0$ be an integer. Let $\theta$ be a real constant with $40<\theta<\pi/$ 4, and let $\mathcal{C}$ be a collection of cones as in Section 4.1.
For any cone $C$ in $\mathcal{C}$, we fix an arbitrary ray $\ell_{C}$ that emanates from the origin and is completely contained in $C$. For any point $p$ in $S$ and any cone $C$ in $\mathcal{C}$, let $C+p$ denote the cone with apex $p$ obtained by translating $C$, i.e., $C+p=\{x+p:x\in C\}$. Similarly, we denote by $\ell_{C}+p$ the ray emanating from $p$ that is parallel to $\ell_{C}$. We denote by $S_{p,C}$ the set of all points in $S\setminus\{p\}$ that are in the cone $C+p$, i.e., $S_{p,C}=(C+p)\cap(S\setminus\{p\})$.
We are now ready to define the variants of the Yao- and $\Theta$-graphs; refer to Figure 8.
Figure 8: The left figure illustrates Yao$(\theta,2f+1)$, where $f=2$. The graph has $52f+1=$ 5 (blue) edges in the cone $C+p$ that are incident on the point $p$; the endpoints of these edges are the five points in this cone that are closest to $p$. The right figure illustrates $\Theta(\theta,2f+1)$, where $f=2$ and $\ell_{C}+p$ is the bisector of the cone $C+p$. The graph has $52f+1=$ 5 (blue) edges in the cone $C+p$ that are incident on the point $p$; the endpoints of these edges are the five points whose orthogonal projections onto the ray $\ell_{C}+p$ are closest to $p$.
  1. 1.
    The graph Yao$(\theta,2f+1)$ has vertex set $S$. For each point $p$ in $S$ and for each cone $C$ in $\mathcal{C}$, we add an edge between $p$ and each of the $\min(2f+1,|S_{p,C}|)$ points in $S_{p,C}$ that are closest (with respect to the Euclidean distance) to $p$.
  2. 2.
    The graph $\Theta(\theta,2f+1)$ also has vertex set $S$. For each point $p$ in $S$ and for each cone $C$ in $\mathcal{C}$, consider again the set $S_{p,C}$. We add an edge between $p$ and each of the $\min(2f+1,|S_{p,C}|)$ points in $S_{p,C}$, whose orthogonal projections onto the ray $\ell_{C}+p$ are closest (with respect to the Euclidean distance) to $p$.
It is clear that the number of edges in each of these graphs is at most $|\mathcal{C}|\cdot(2f+1)n=O(fn)$.
4.3 Analysis
In this subsection, we will prove that each of the graphs Yao$(\theta,2f+1)$ and $\Theta(\theta,2f+1)$ is an $f$-faulty-degree $t$-spanner for $S$, where $t$ only depends on the angle $\theta$.
We remark that Lukovszki [15] used $\Theta(\theta,f+1)$ for the case when at most $f$ edges can be removed. Our proof is similar.
Our proof will use the following geometric result, which is illustrated in Figure 9.
Figure 9: Illustrating the first case in Lemma 9.
Lemma 9
Let $p$ and $q$ be two distinct points in $S$, and let $C$ be a cone in $\mathcal{C}$ such that $q\in C+p$.
  1. 1.
    Let $\ell$ be an arbitrary ray that emanates from $p$ and is completely contained in $C+p$. Let $r$ be a point in $S_{p,C}$ such that the orthogonal projection of $r$ onto $\ell$ is at least as close to $p$ as the orthogonal projection of $q$ onto $\ell$. Then,
    $|rq|\leq|pq|-(\cos\theta-\sin\theta)|pr|.$
  2. 2.
    Let $r$ be a point in $S_{p,C}$ such that $|pr|\leq|pq|$. Then,
    $|rq|\leq|pq|-(\cos\theta-\sin\theta)|pr|.$
Proof. The proof of the first claim can be found in [17, Lemma 4.1.4]. The second claim follows from the first claim, by taking for $\ell$ the ray that emanates from $p$ and goes through $q$.
In the rest of this subsection, we will prove that each of the graphs Yao$(\theta,2f+1)$ and $\Theta(\theta,2f+1)$ is an $f$-faulty-degree $t$-spanner for $S$, where $t=1/(\cos\theta-\sin\theta)$. For both graphs, the proof will be by induction. For Yao$(\theta,2f+1)$, the second claim in Lemma 9 is used, whereas for $\Theta(\theta,2f+1)$, the first claim is used. Since both proofs are almost identical, we only give the proof for $\Theta(\theta,2f+1)$.
For simplicity, we denote the graph $\Theta(\theta,2f+1)$ by $\Theta$. Let $F$ be an arbitrary subset of the edge set of $\Theta$ and assume that the maximum degree of the graph $(S,F)$ is at most $f$.
Lemma 10
Let $\{p,q\}$ be an edge in $K_{S}\setminus F$ that is not in $\Theta$. Then, there is a point $r$ in $S_{p,C}$ such that
  • $r\neq q$,
  • the orthogonal projection of $r$ onto the ray $\ell_{C}+p$ is at least as close to $p$ as the orthogonal projection of $q$ onto $\ell_{C}+p$,
  • $\{p,r\}$ is an edge in $\Theta\setminus F$, and
  • $\{r,q\}$ is an edge in $K_{S}\setminus F$.
Proof. Since $\{p,q\}$ is not an edge in $\Theta$, this graph contains $12f+$ 1 edges $\{p,r\}$, where $r$ is in $S_{p,C}$ and $r\neq q$. For each such edge $\{p,r\}$, the orthogonal projection of $r$ onto the ray $\ell_{C}+p$ is at least as close to $p$ as the orthogonal projection of $q$ onto $\ell_{C}+p$. Since the degree of $p$ in $F$ is at most $f$, at least $f+1$ of these edges $\{p,r\}$ are in $\Theta\setminus F$. Observe that for each such edge $\{p,r\}$, the edge $\{r,q\}$ is in $K_{S}$. Since the degree of $q$ in $F$ is at most $f$, for at least one of these $f+1$ edges $\{p,r\}$ that are in $\Theta\setminus F$, the edge $\{r,q\}$ is in $K_{S}\setminus F$.
Lemma 11
For every edge $\{p,q\}$ in $K_{s}\setminus F$, $\delta_{\Theta\setminus F}(p,q)\leq t\cdot|pq|$, where $t=1/(\cos\theta-\sin\theta)$.
Proof. Let $C$ be a cone in $\mathcal{C}$ such that $q\in C+p$. We will prove the lemma by induction on the rank of the Euclidean distance $|pq|$ in the set of all edge lengths in $K_{S}\setminus F$.
For the base case, assume that $\{p,q\}$ is a shortest edge in $K_{s}\setminus F$. We claim that this edge is in $\Theta$. If this is true, then it is in $\Theta\setminus F$ as well and $\delta_{\Theta\setminus F}(p,q)=|pq|\leq t\cdot|pq|$. Assume that $\{p,q\}$ is not in $\Theta$. Let $r$ be a point in $S_{p,C}$ as in Lemma 10. It follows from the first claim in Lemma 9, with $\ell=\ell_{C}+p$, that $|rq|<|pq|$, which contradicts the assumption that $\{p,q\}$ is a shortest edge in $K_{s}\setminus F$.
From now on, assume that $\{p,q\}$ is not a shortest edge in $K_{s}\setminus F$. If this edge is in $\Theta$, then it is in $\Theta\setminus F$ and the lemma holds. Otherwise, let $r$ be a point in $S_{p,C}$ as in Lemma 10. As above, by Lemma 9, we have $|rq|<|pq|$. Thus, by induction,
$\delta_{\Theta\setminus F}(r,q)\leq t\cdot|rq|.$
It follows that
$\delta_{\Theta\setminus F}(p,q)\leq|pr|+\delta_{\Theta\setminus F}(r,q)\leq|pr|+t\cdot|rq|.$
Using the first claim in Lemma 9, with $\ell=\ell_{C}+p$, a straightforward calculation shows that
$\delta_{\Theta\setminus F}(p,q)\leq t\cdot|pq|.$
Theorem 4
Let $S$ be a set of $n$ points in $\mathbb{R}^{d}$, where $d$ is a constant, let $\varepsilon>0$ be a real constant, let $\theta$ be a constant with $40<\theta<\pi/$ 4, such that $1/(\cos\theta-\sin\theta)\leq 1+\varepsilon$, and let $f\geq 0$ be an integer. Each of the graphs Yao$(\theta,2f+1)$ and $\Theta(\theta,2f+1)$ is an $f$-faulty-degree $(1+\varepsilon)$-spanner for $S$, having $O(fn)$ edges.
Proof. We only give the proof for $\Theta=\Theta(\theta,2f+1)$. Set $t=1/(\cos\theta-\sin\theta)$. Let $F$ be an arbitrary subset of the edge set of $\Theta$, such that the maximum degree of the graph $(S,F)$ is at most $f$. Let $p$ and $q$ be two distinct points in $S$. If $\{p,q\}$ is an edge in $K_{S}\setminus F$, then, by Lemma 11,
$\delta_{\Theta\setminus F}(p,q)\leq t\cdot|pq|=t\cdot\delta_{K_{s}\setminus F}(p,q)\leq(1+\varepsilon)\cdot\delta_{K_{s}\setminus F}(p,q).$
If $\{p,q\}$ is not an edge in $K_{S}\setminus F$, then we consider a shortest path between $p$ and $q$ in $K_{S}\setminus F$, and apply Lemma 11 to each edge on this path.