### v2

# Metric and Geometric Spanners that are Resilient to Degree-Bounded Edge Faults

Abstract
Let $H$ be an edge-weighted graph, and let $G$ be a subgraph of $H$. We say that $G$ is an $f$-fault-tolerant $t$-spanner for $H$, if the following is true for any subset $F$ of at most $f$ edges of $G$: For any two vertices $p$ and $q$, the shortest-path distance between $p$ and $q$ in the graph $G\setminus F$ is at most $t$ times the shortest-path distance between $p$ and $q$ in the graph $H\setminus F$. The value of $t$ is called the stretch factor. Numerous results are known for constructing such spanners, both for general graphs $H$ and for complete graphs $H$ whose vertex set is a set of points in $\mathbb{R}^{d}$ or in a metric space of bounded doubling dimension.
Bodwin, Haeupler, and Parter (SODA 2024) generalized this notion to the case when $F$ can be any set of edges in $G$, as long as the maximum degree of $F$ is at most $f$. For example, if $f=1$, then $F$ can be a perfect matching in $G$. They gave constructions for general graphs $H$.
We present three new results for constructing spanners that are resilient to edge faults of maximum degree $f$.
We first consider the case when $H$ is a complete graph whose vertex set is an arbitrary metric space. We show that if this metric space contains a $t$-spanner with $m$ edges, then it also contains a graph $G$ with at most (4f1)m(4f-1)m edges, that is resilient to edge faults of maximum degree $f$ and has stretch factor $O(ft)$. We also give an example for which our construction has stretch factor $\Omega(ft)$. If $f=1$, the graph $G$ has at most $3m$ edges and its stretch factor is at most $3t$. For this case, we give an example for which $G$ has stretch factor arbitrarily close to $3t$.
Next, we consider the case when $H$ is a complete graph whose vertex set is a metric space that admits a well-separated pair decomposition. We show that, if the metric space has such a decomposition of size $m$, then it contains a graph with at most $(2f+1)^{2}m$ edges, that is resilient to edge faults of maximum degree $f$ and has stretch factor at most $1+\varepsilon$, for any given $\varepsilon>0$. For example, if the vertex set is a set of $n$ points in $\mathbb{R}^{d}$ ($d$ being a constant) or a set of $n$ points in a metric space of bounded doubling dimension, then the spanner has $O(f^{2}n)$ edges.
Finally, for the case when $H$ is a complete graph on $n$ points in $\mathbb{R}^{d}$, we show how natural variants of the Yao- and $\Theta$-graphs lead to graphs with $O(fn)$ edges, that are resilient to edge faults of maximum degree $f$ and have stretch factor at most $1+\varepsilon$, for any given $\varepsilon>0$.
1 Introduction
Let $H$ be an undirected “host graph”, in which each edge has a weight. For a real number $t\geq 1$, a subgraph $G$ of $H$ is called a $t$-spanner for $H$, if for any two vertices $p$ and $q$ of $H$,
$\delta_{G}(p,q)\leq t\cdot\delta_{H}(p,q),$
where, for any edge-weighted graph $X$, $\delta_{X}(p,q)$ denotes the length of a shortest path between $p$ and $q$ in $X$. The smallest such value for $t$ is called the stretch factor of $G$.
There is a large amount of literature on the problem of designing algorithms that compute a $t$-spanner for a graph $H$, where the goal is to obtain the best trade-offs between the stretch factor $t$ and the number of edges in $G$. For general edge-weighted graphs $H$, we refer the reader to the survey paper Ahmed et al. [1].
In the metric case, the host graph $H$ is a complete graph $K_{S}$ induced by a finite metric space $(S,|\cdot|)$: The vertex set of $K_{S}$ is $S$ and for any two distinct points $p$ and $q$ in $S$, $K_{S}$ contains the edge $\{p,q\}$ with weight $|pq|$. In this case, a graph $G$ with vertex set $S$ is a $t$-spanner for $S$, if for any two points $p$ and $q$ of $S$,
$\delta_{G}(p,q)\leq t\cdot|pq|.$
For several classes of metric spaces, constructions are known that compute, for any constant $t>1$, a $t$-spanner with $O(n)$ edges, where $n$ is the number of points in $S$. For example, this is the case for Euclidean spaces, where $S$ is a finite set of points in $\mathbb{R}^{d}$ (where the dimension $d$ is a constant) and the distance between two points is their Euclidean distance. This was first shown by Keil [14]. Many other constructions can be found in Narasimhan and Smid [19]. Another example is the case when the metric space $(S,|\cdot|)$ has bounded doubling dimension. For this case, Har-Peled and Mendel [12] were the first to show how to compute a $t$-spanner with $O(n)$ edges. Other results can be found in Filtser and Solomon [9].
1.1 Fault-Tolerant Spanners
Spanners were generalized by Levcopoulos et al. [15, 16] to fault-tolerant spanners. Let $f\geq 1$ be an integer. A subgraph $G$ of the host graph $H$ is an $f$-fault-tolerant $t$-spanner for $H$, if the following holds for any subset $F$ of at most $f$ edges of $G$: For any two vertices $p$ and $q$ of $H$,
$\delta_{G\setminus F}(p,q)\leq t\cdot\delta_{H\setminus F}(p,q),$ (1)
where, for any edge-weighted graph $X=(V_{X},E_{X})$, $X\setminus F$ denotes the graph with vertex set $V_{X}$ and edge set $E_{X}\setminus F$. Thus, $\delta_{G\setminus F}(p,q)$ denotes the length of a shortest path between $p$ and $q$ in the graph obtained by removing all edges in $F$ from $G$. We remark that there are also versions of fault-tolerant spanners in which up to $f$ vertices and edges can be removed.
Again, there is a large amount of literature on the problem of constructing fault-tolerant spanners. For the case of general edge-weighted graphs, the first results were given by Chechik et al. [7]. See Bodwin et al. [3] for more recent results. For Euclidean spaces and metric spaces of bounded doubling dimension, we refer the reader to [6, 13, 17, 18, 19, 20].
Bodwin, Haeupler, and Parter [4] introduced a more general notion of fault-tolerant spanners. A subgraph $G=(V_{H},E)$ of the host graph $H=(V_{H},E_{H})$ is an $f$-faulty-degree $t$-spanner for $H$, if for any subset $F$ of $E$, for which the maximum degree of the graph $(V_{H},F)$ is at most $f$, and for any two vertices $p$ and $q$ in $V_{H}$, (1) holds.
For example, in a $1$-faulty-degree $t$-spanner, we are allowed to remove a perfect matching that has $|V_{H}|/2$ edges, whereas in a $1$-fault-tolerant $t$-spanner, we are only allowed to remove one edge. This shows that an $f$-faulty-degree spanner is a much stronger condition.
Bodwin, Haeupler, and Parter [4] proved the following result. Let $H$ be any edge-weighted graph with $n$ vertices. For any positive integers $f$ and $k$, there exists an $f$-faulty-degree (2k1)(2k-1)-spanner for $H$ that has $O(k^{k}\cdot f^{1-1/k}\cdot n^{1+1/k})$ edges. Notice that there is a large dependance on $k$. In particular, the upper bound on the number of edges is subquadratic in $n$ only if $k=o(\log n/\log\log n)$. [^1] Recall that $k^{k}=n$ solves to $k\sim\log n/\log\log n$.
1.2 Our Results
We consider faulty-degree spanners for the case when the host graph is the complete graph $K_{S}$ induced by a finite metric space $(S,|\cdot|)$. For completeness, we state the definition of a faulty-degree spanner for this case.
Definition 1
Let $(S,|\cdot|)$ be a finite metric space, let $f\geq 0$ be an integer, let $t\geq 1$ be a real number, and let $G=(S,E)$ be an undirected graph, in which each edge $\{p,q\}$ has weight $|pq|$. The graph $G$ is called an $f$-faulty-degree $t$-spanner for $S$, if the following holds for any subset $F$ of $E$, for which the maximum degree of the graph $(S,F)$ is at most $f$: For any two points $p$ and $q$ in $S$,
$\delta_{G\setminus F}(p,q)\leq t\cdot\delta_{K_{s}\setminus F}(p,q).$
In Section 2, we consider arbitrary metric spaces $(S,|\cdot|)$. We will show the following: Assume that $G^{\prime}=(S,E^{\prime})$ is an arbitrary $t$-spanner for $S$. Let $n$ denote the number of points in $S$, and let $m$ denote the number of edges in $E^{\prime}$. For any integer $f\geq 1$, we can convert $G^{\prime}$ to a graph $G=(S,E)$ with at most (4f1)m(4f-1)m edges, such that $G$ is an $f$-faulty-degree $t^{\prime}$-spanner for $S$, where $t^{\prime}$ is a function of $f$ and $t$. The construction is valid for any $f$ with $1\leq f\leq(n-1)/2$. In Section 2.1, we consider the case when $f=1$ and prove that $t^{\prime}\leq 3t$. We also give an example for which $t^{\prime}$ is arbitrarily close to $3t$. In Section 2.2, we consider the general case $f\geq 1$ and prove that $t^{\prime}\leq(8f+2)t$. For $f\geq 3$, we give an example for which $t^{\prime}=\Omega(ft)$.
In Section 3, we assume that $S$ is a set of $n$ points in a metric space that admits a well-separated pair decomposition (WSPD). Let $m$ denote the size of such a WSPD. We will show that, for any integer $f\geq 1$ and any real constant $\varepsilon>0$, there exists an $f$-faulty-degree $(1+\varepsilon)$-spanner for $S$ that has at most $(2f+1)^{2}m$ edges. We remark that our spanner is obtained by modifying the construction in Levcopoulos et al. [16] for standard fault-tolerant spanners. Examples of metric spaces for which $m=O(n)$ are Euclidean spaces of constant dimension (see Callahan and Kosaraju [5]) and metric spaces of bounded doubling dimension (see Har-Peled and Mendel [12]).
In Section 4, we will show that in Euclidean space, we can do even better: Assume that $S$ is a set of $n$ points in $\mathbb{R}^{d}$, where the dimension $d$ is a constant, and the distance $|pq|$ between two points $p$ and $q$ is their Euclidean distance. For this case, we will define two graphs with vertex set $S$ that are variants of the Yao-graph (see Flinchbaugh and Jones [10] and Yao [21]) and the $\Theta$-graph (see Clarkson [8] and Keil [14]) and that depend on the maximum degree bound $f$ of any set $F$ of faulty edges. We will prove that, for any real constant $\varepsilon>0$, each of these graphs is an $f$-faulty-degree $(1+\varepsilon)$-spanner for $S$ with $O(fn)$ edges. In particular, the stretch factor does not depend on $f$. Observe that this is optimal, up to constant factors, because every vertex in an $f$-faulty-degree spanner must have degree at least $f$. We remark that our graphs are variations of the standard fault-tolerant spanners obtained by Lukovszki [17].
We conclude this section by introducing notation that will be used throughout the paper. Let $P=(p_{1},p_{2},\ldots,p_{k})$ be a sequence of points in a finite metric space $(S,|\cdot|)$. The length of this sequence is denoted by $|P|$ and is equal to $\sum_{i=1}^{k-1}|p_{i}p_{i+1}|$. If all points are pairwise distinct, then $P$ is a path in $K_{S}$. If the sequence may have repeated points, then $P$ is a walk in $K_{S}$.
2 General Metric Spaces
In this section, $(S,|\cdot|)$ denotes an arbitrary finite metric space. Let $n$ be the number of points in $S$ and let $t\geq 1$ be a real number. We assume that we are given a $t$-spanner $G^{\prime}=(S,E^{\prime})$ for $S$ and we denote the number of edges of $G^{\prime}$ by $m$. For any integer $f$ with $1\leq f\leq(n-1)/2$, we will show how to convert $G^{\prime}$ to a graph $G=(S,E)$ that has at most (4f1)m(4f-1)m edges and is an $f$-faulty-degree $t^{\prime}$-spanner for $S$, where $t^{\prime}$ is a function of $f$ and $t$.
The graph $G=(S,E)$ is obtained as follows (refer to Figure 1):
Figure 1: Illustrating the edges that are added to $E$ because of the edge $\{a,b\}$ in $E^{\prime}$. In this figure, $f=3$.
  • The edge set $E$ contains all edges of $E^{\prime}$.
  • For each edge $\{a,b\}$ in $E^{\prime}$, let $c_{1},c_{2},\ldots,c_{n-2}$ be the sequence of points of $S\setminus\{a,b\}$, in non-decreasing order of the values $|ac_{i}|+|c_{i}b|$; ties are broken arbitrarily. Let $C_{ab}$ denote the set consisting of the first 2f12f-1 points in this sequence, i.e., $C_{ab}=\{c_{1},c_{2},\ldots,c_{2f-1}\}$. The edge set $E$ contains the 4f24f-2 edges $\{a,c_{i}\}$ and $\{c_{i},b\}$, for $i=1,2,\ldots,2f-1$.
Observe that the number of edges in $E$ is at most m+(4f2)m=(4f1)mm+(4f-2)m=(4f-1)m.
In the rest of this section, we will prove that $G$ is an $f$-faulty-degree $t^{\prime}$-spanner for $S$. In Section 2.1, we consider the case when $f=1$, in which case $t^{\prime}\leq 3t$. In Section 2.2, we consider the general case when $f\geq 1$ and prove that $t^{\prime}\leq(8f+2)t$.
2.1 Faulty Matchings: $f=1$
Consider the graph $G=(S,E)$ defined above for the case when $f=1$. Any subset $F$ of $E$ such that the maximum degree of the graph $(S,F)$ is at most $1$ is a matching in $G$; this matching may not be a perfect matching. In other words, $G\setminus F$ is obtained by removing this matching from $G$. We will prove the following result.
Theorem 1
Let $(S,|\cdot|)$ be a finite metric space, let $t\geq 1$ be a real number, and let $G^{\prime}=(S,E^{\prime})$ be a $t$-spanner for $S$ with $m$ edges. Then the graph $G=(S,E)$ is a $1$-faulty-degree $(3t)$-spanner for $S$ that has at most $3m$ edges.
We have already seen that the number of edges in $E$ is at most $3m$. It remains to prove that $G$ is a $1$-faulty-degree $(3t)$-spanner for $S$.
Let $F$ be an arbitrary matching in $G$, and let $p$ and $q$ be two points in $S$. We have to show that
$\delta_{G\setminus F}(p,q)\leq 3t\cdot\delta_{K_{s}\setminus F}(p,q).$
Let $P^{\prime}$ be a shortest path between $p$ and $q$ in the graph $G^{\prime}$. Since $G^{\prime}$ is a $t$-spanner for $S$, the length $|P^{\prime}|$ of $P^{\prime}$ is at most $t|pq|$.
Let $\{a_{1},b_{1}\},\{a_{2},b_{2}\},\ldots,\{a_{k},b_{k}\}$ be the edges on $P^{\prime}$ that are in $F$. We may assume that, if we follow $P^{\prime}$ from $p$ to $q$, the endpoints of these edges are visited in the order $a_{1},b_{1},a_{2},b_{2},\ldots,a_{k},b_{k}$; refer to Figure 2. Since $F$ is a matching, $b_{1}\neq a_{2},b_{2}\neq a_{3},\ldots,b_{k-1}\neq a_{k}$.
Figure 2: $P^{\prime}$ is a shortest path between $p$ and $q$ in the graph $G^{\prime}$. The red edges $\{a_{i},b_{i}\}$, $1\leq i\leq k$, are in the matching $F$. $P$ (in green) is a walk between $p$ and $q$ in the graph $G\setminus F$. Note that any vertex $c_{i}$ may be on the path $P^{\prime}$.
For each $i$ with $1\leq i\leq k$, let $c_{i}$ be the point in $S\setminus\{a_{i},b_{i}\}$ that minimizes $|a_{i}c_{i}|+|c_{i}b_{i}|$ and that was chosen when defining the graph $G$. Since $F$ is a matching, all edges $\{a_{i},c_{i}\}$ and $\{c_{i},b_{i}\}$, for $1\leq i\leq k$, are in $G\setminus F$.
We define $P$ to be the walk (i.e., possibly non-simple path) that
  • starts at $p$ and follows $P^{\prime}$ to $a_{1}$,
  • for $i=1,2,\ldots,k-1$, takes the two edges $\{a_{i},c_{i}\}$ and $\{c_{i},b_{i}\}$, and then follows $P^{\prime}$ from $b_{i}$ to $a_{i+1}$,
  • takes the two edges $\{a_{k},c_{k}\}$ and $\{c_{k},b_{k}\}$, and then follows $P^{\prime}$ from $b_{k}$ to $q$.
Refer to Figure 2. Observe that $P$ is a walk between $p$ and $q$ in the graph $G\setminus F$.
If $k=0$, then $P=P^{\prime}$ and we have
$\delta_{G\setminus F}(p,q)=|P^{\prime}|\leq t|pq|=t\cdot\delta_{K_{S}\setminus F}(p,q)\leq 3t\cdot\delta_{K_{S}\setminus F}(p,q).$
Assume that $k=1$ and $P^{\prime}$ consists of just the edge $\{p,q\}$. Then $p=a_{1}$, $q=b_{1}$, and the length of $P$ is equal to $|a_{1}c_{1}|+|c_{1}b_{1}|$. By the definition of $c_{1}$, we have $\delta_{K_{S}\setminus F}(p,q)=|a_{1}c_{1}|+|c_{1}b_{1}|$ and, thus,
$\delta_{G\setminus F}(p,q)=\delta_{K_{S}\setminus F}(p,q)\leq 3t\cdot\delta_{K_{S}\setminus F}(p,q).$
It remains to consider the case covered by the following lemma.
Lemma 1
Assume that $k\geq 1$ and $P^{\prime}$ has at least two edges. Then the graph $G\setminus F$ contains a walk $Q$ between $p$ and $q$ whose length is at most $3|P^{\prime}|$.
This lemma will imply that
$\delta_{G\setminus F}(p,q)\leq|Q|\leq 3|P^{\prime}|\leq 3t|pq|\leq 3t\cdot\delta_{K_{s}\setminus F}(p,q).$
2.1.1 Proof of Lemma 1
For any two vertices $x$ and $y$ of $P^{\prime}$, we denote by $P^{\prime}_{xy}$ the subpath of $P^{\prime}$ between $x$ and $y$. Using this notation, the length $|P|$ of $P$ satisfies
$|P|=|P^{\prime}_{pa_{1}}|+\sum_{i=1}^{k-1}|P^{\prime}_{b_{i}a_{i+1}}|+|P^{\prime}_{b_{k}q}|+\sum_{i=1}^{k}\left(|a_{i}c_{i}|+|c_{i}b_{i}|\right).$ (2)
To prove Lemma 1, we consider four cases. In order to verify that these cases cover all possibilities, we describe them first:
  • Case 1: At least one of the two edges of $P^{\prime}$ that are incident on $p$ and $q$ is not in $F$.
  • In the remaining three cases, both edges of $P^{\prime}$ that are incident on $p$ or $q$ are in $F$.
  • Assume that $k=2$. Since $F$ is a matching in $G$, and $G^{\prime}$ is a subgraph of $G$, $P^{\prime}$ has at least three edges.
  • Case 2: $k=2$ and $P^{\prime}$ consists of three edges.
  • Case 3: $k=2$ and $P^{\prime}$ consists of at least four edges.
  • Case 4: $k\geq 3$.
Case 1: The edge of $P^{\prime}$ that is incident on $p$ is not in $F$ or the edge of $P^{\prime}$ that is incident on $q$ is not in $F$.
By symmetry, we may assume that the edge of $P^{\prime}$ that is incident on $p$ is not in $F$. Since $p\not\in\{a_{1},b_{1}\}$, our choice of $c_{1}$ implies that
$|a_{1}c_{1}|+|c_{1}b_{1}|\leq|a_{1}p|+|pb_{1}|.$
For each $i$ with $2\leq i\leq k$, since $b_{i-1}\not\in\{a_{i},b_{i}\}$, our choice of $c_{i}$ implies that
$|a_{i}c_{i}|+|c_{i}b_{i}|\leq|a_{i}b_{i-1}|+|b_{i-1}b_{i}|.$
It follows that
$\sum_{i=1}^{k}\left(|a_{i}c_{i}|+|c_{i}b_{i}|\right)$ $\leq$ $\left(|a_{1}p|+|pb_{1}|\right)+\sum_{i=2}^{k}\left(|a_{i}b_{i-1}|+|b_{i-1}b_{i}|\right)$
$\leq$ $|P^{\prime}_{pa_{1}}|+|P^{\prime}_{pb_{1}}|+\sum_{i=2}^{k}\left(|P^{\prime}_{b_{i-1}a_{i}}|+|P^{\prime}_{b_{i-1}b_{i}}|\right)$
$\leq$ $2|P^{\prime}|.$
Using this, it follows from (2) that $|P|\leq 3|P^{\prime}|$. Hence, Lemma 1 holds for $Q=P$.
Case 2: $k=2$, the path $P^{\prime}$ consists of three edges, the edge of $P^{\prime}$ that is incident on $p$ is in $F$, and the edge of $P^{\prime}$ that is incident on $q$ is in $F$.
In this case, $p=a_{1}$, $q=b_{2}$, and the path $P$ consists of the five edges $\{a_{1},c_{1}\}$, $\{c_{1},b_{1}\}$, $\{b_{1},a_{2}\}$, $\{a_{2},c_{2}\}$, and $\{c_{2},b_{2}\}$. Observe that $\{b_{1},a_{2}\}$ is not in $F$.
Let $c$ be the point in $S\setminus\{b_{1},a_{2}\}$ that minimizes $|b_{1}c|+|ca_{2}|$ and that was chosen when defining the graph $G$.
Case 2.1: $c\not\in\{a_{1},b_{2}\}$.
Since $a_{1}\not\in\{b_{1},a_{2}\}$, our choice of $c$ implies that
$|b_{1}c|+|ca_{2}|\leq|b_{1}a_{1}|+|a_{1}a_{2}|\leq|b_{1}a_{1}|+|a_{1}b_{1}|+|b_{1}a_{2}|.$
Similarly, since $b_{2}\not\in\{b_{1},a_{2}\}$,
$|b_{1}c|+|ca_{2}|\leq|b_{1}b_{2}|+|b_{2}a_{2}|\leq|b_{1}a_{2}|+|a_{2}b_{2}|+|b_{2}a_{2}|.$
By adding these two inequalities, we get
$2\left(|b_{1}c|+|ca_{2}|\right)\leq 2|a_{1}b_{1}|+2|b_{1}a_{2}|+2|a_{2}b_{2}|=2|P^{\prime}|,$
which gives
$|b_{1}c|+|ca_{2}|\leq|P^{\prime}|.$ (3)
Since $c\not\in\{a_{1},b_{1}\}$, our choice of $c_{1}$ implies that
$|a_{1}c_{1}|+|c_{1}b_{1}|\leq|a_{1}c|+|cb_{1}|.$ (4)
Since $c\not\in\{a_{2},b_{2}\}$, our choice of $c_{2}$ implies that
$|a_{2}c_{2}|+|c_{2}b_{2}|\leq|a_{2}c|+|cb_{2}|.$ (5)
Using (4) and (5), we get
$|P|$ $=$ $|a_{1}c_{1}|+|c_{1}b_{1}|+|b_{1}a_{2}|+|a_{2}c_{2}|+|c_{2}b_{2}|$
$\leq$ $|a_{1}c|+|cb_{1}|+|b_{1}a_{2}|+|a_{2}c|+|cb_{2}|.$
Using (3), we get
$|P|$ $\leq$ $|a_{1}c|+|b_{1}a_{2}|+|cb_{2}|+|P^{\prime}|$
$\leq$ $\left(|a_{1}b_{1}|+|b_{1}c|\right)+|b_{1}a_{2}|+\left(|ca_{2}|+|a_{2}b_{2}|\right)+|P^{\prime}|$
$=$ $|b_{1}c|+|ca_{2}|+2|P^{\prime}|.$
One more application of (3) gives
$|P|\leq 3|P^{\prime}|.$
Hence, Lemma 1 holds for $Q=P$.
Case 2.2: $c\in\{a_{1},b_{2}\}$.
By symmetry, we may assume that $c=a_{1}$. By the definition of the graph $G$, $\{c,a_{2}\}=\{p,a_{2}\}$ is an edge in $G$. Since $F$ is a matching, this edge is in the graph $G\setminus F$. Define the path $Q=(p,a_{2},c_{2},b_{2}=q)$. This path is in $G\setminus F$.
Since $c\not\in\{a_{2},b_{2}\}$, our choice of $c_{2}$ implies that
$|a_{2}c_{2}|+|c_{2}b_{2}|\leq|a_{2}c|+|cb_{2}|,$
i.e.,
$|a_{2}c_{2}|+|c_{2}q|\leq|a_{1}a_{2}|+|a_{1}b_{2}|.$
The length of the path $Q$ satisfies
$|Q|=|pa_{2}|+|a_{2}c_{2}|+|c_{2}q|\leq|pa_{2}|+|a_{1}a_{2}|+|a_{1}b_{2}|.$
By the triangle inequality, we have $|Q|\leq 3|P^{\prime}|$. Hence, Lemma 1 holds for $Q$.
Case 3: $k=2$, the path $P^{\prime}$ consists of at least four edges, the edge of $P^{\prime}$ that is incident on $p$ is in $F$, and the edge of $P^{\prime}$ that is incident on $q$ is in $F$.
In this case, $p=a_{1}$ and $q=b_{2}$. Let $z$ be the third point on $P^{\prime}$, when traversing this path from $p$ to $q$. Then, the edge $\{b_{1},z\}$ is on $P^{\prime}$ and $z\neq a_{1}$.
Observe that $P^{\prime}_{pz}$ is a shortest path between $p$ and $z$ in $G^{\prime}$ and this path satisfies the condition of Case 2. Therefore, the graph $G\setminus F$ contains a walk $Q_{1}$ between $p$ and $z$ whose length is at most $3|P^{\prime}_{pz}|$.
Similarly, $P^{\prime}_{zq}$ is a shortest path between $z$ and $q$ in $G^{\prime}$ and this path satisfies the condition of Case 1. Thus, the graph $G\setminus F$ contains a walk $Q_{2}$ between $z$ and $q$ whose length is at most $3|P^{\prime}_{zq}|$.
Let $Q$ be the concatenation of $Q_{1}$ and $Q_{2}$. Then $Q$ is a walk between $p$ and $q$ in $G\setminus F$. The length of this walk satisfies
$|Q|=|Q_{1}|+|Q_{2}|\leq 3|P^{\prime}_{pz}|+3|P^{\prime}_{zq}|=3|P^{\prime}|.$
Thus, Lemma 1 holds for this walk $Q$.
Case 4: $k\geq 3$, the edge of $P^{\prime}$ that is incident on $p$ is in $F$, and the edge of $P^{\prime}$ that is incident on $q$ is in $F$.
In this case, $p=a_{1}$ and $q=b_{k}$. Observe that $P^{\prime}_{pb_{2}}$ is a shortest path between $p$ and $b_{2}$ in $G^{\prime}$ and this path satisfies one of the conditions of Cases 3 and 4. Therefore, the graph $G\setminus F$ contains a walk $Q_{1}$ between $p$ and $b_{2}$ whose length is at most $3|P^{\prime}_{pb_{2}}|$. Similarly, $P^{\prime}_{b_{2}q}$ is a shortest path between $b_{2}$ and $q$ in $G^{\prime}$ and this path satisfies the condition of Case 1. Thus, the graph $G\setminus F$ contains a walk between $b_{2}$ and $q$ whose length is at most $3|P^{\prime}_{b_{2}q}|$. Let $Q$ be the concatenation of $Q_{1}$ and $Q_{2}$. Then $Q$ is a walk between $p$ and $q$ in $G\setminus F$. The length of this walk satisfies
$|Q|=|Q_{1}|+|Q_{2}|\leq 3|P^{\prime}_{pb_{2}}|+3|P^{\prime}_{b_{2}q}|=3|P^{\prime}|.$
Thus, Lemma 1 holds for this walk $Q$.
This completes the proof of Lemma 1 and, therefore, of Theorem 1.
2.1.2 Lower Bound
We show that the upper bound of $3t$ on the stretch factor of the graph $G$ in Theorem 1 is tight. Let $\varepsilon$ be a real number with $0<\varepsilon\ll 1$, and define $\varepsilon^{\prime}=\varepsilon/(4-2\varepsilon)$. Consider the one-dimensional point set $S=\{p,a,b,q\}$, as indicated below. The graph $G^{\prime}$ connects the points in the left-to-right order and is a $1$-spanner for $S$. The graph $G$ contains $G^{\prime}$ and the two edges $\{p,b\}$ and $\{a,q\}$.
Consider the matching $F=\{\{p,a\},\{b,q\}\}$, as indicated in red. We have $\delta_{G\setminus F}(p,q)=3+2\varepsilon^{\prime}$ and $\delta_{K_{S}\setminus F}(p,q)=1+2\varepsilon^{\prime}$. Thus,
$\frac{\delta_{G\setminus F}(p,q)}{\delta_{K_{S}\setminus F}(p,q)}=\frac{3+2\varepsilon^{\prime}}{1+2\varepsilon^{\prime}}=3-\varepsilon,$
i.e., $G^{\prime}$ is a $1$-faulty-degree $(3-\varepsilon)$-spanner for $S$.
2.2 The General Case: $f\geq 1$
Consider the graph $G=(S,E)$ defined in the beginning of Section 2 for the general case when $f\geq 1$. We will prove the following result.
Theorem 2
Let $(S,|\cdot|)$ be a finite metric space, let $n$ be the number of points in $S$, let $t\geq 1$ be a real number, and let $G^{\prime}=(S,E^{\prime})$ be a $t$-spanner for $S$ with $m$ edges. Then, for any integer $f$ with $1\leq f\leq(n-1)/2$, the graph $G=(S,E)$ is an $f$-faulty-degree $((8f+2)t)$-spanner for $S$ that has at most (4f1)m(4f-1)m edges.
We have already seen that the number of edges in $E$ is at most (4f1)m(4f-1)m. It remains to prove that $G$ is an $f$-faulty-degree $((8f+2)t)$-spanner for $S$.
Let $F$ be an arbitrary subset of the edge set $E$ of $G$, such that the maximum degree of the graph $(S,F)$ is at most $f$. In the rest of this section, we will prove that for any two distinct points $p$ and $q$ in $S$,
$\delta_{G\setminus F}(p,q)\leq(8f+2)t\cdot\delta_{K_{S}\setminus F}(p,q).$ (6)
Since the proof is long and quite intricate, we have broken it up into several preliminary lemmas. The proof of (6) will be given in Lemma 6.
Lemma 2
The graph $K_{S}\setminus F$ is connected.
Proof. We will prove the following claim, which will imply the statement in the lemma: For every partition of $S$ into two non-empty sets $A$ and $B$, the graph $K_{S}\setminus F$ contains an edge between $A$ and $B$.
To prove this claim, we may assume without loss of generality that the size of $B$ is at least the size of $A$. Thus, the size of $B$ is at least $n/2$, which is larger than $f$. Let $p$ be a point in $A$. In $K_{S}$, there are more than $f$ edges of the form $\{p,q\}$, with $q\in B$. At least one these edges is in $K_{S}\setminus F$.
We will repeatedly use the following lemma, which applies to any edge in $G^{\prime}$ that is in the matching $F$.
Lemma 3
Let $\{a,b\}$ be an edge in $G^{\prime}$. If this edge is in $F$, then there exists a point $c$ in $C_{ab}$ such that both $\{a,c\}$ and $\{c,b\}$ are edges in the graph $G\setminus F$.
Proof. Assume that for each of the 2f12f-1 points $c$ in $C_{ab}$, at least one of the two edges $\{a,c\}$ and $\{c,b\}$ is in $F$. Then, in $(S,F)$, the sum of the degrees of $a$ and $b$ is at least 2+(2f1)=2f+12+(2f-1)=2f+1. Therefore, $a$ or $b$ has degree more than $f$ in $(S,F)$, which is a contradiction.
The next lemma states that for any edge $\{a,b\}$ in $G^{\prime}$, the graph $G\setminus F$ contains a shortest path, all of whose edges are “close” to both $a$ and $b$.
Lemma 4
Let $\{a,b\}$ be an edge in $G^{\prime}$. Then there exists a shortest path between $a$ and $b$ in the graph $G\setminus F$, all of whose vertices belong to $C_{ab}\cup\{a,b\}$.
Proof. First assume that $\{a,b\}$ is not an edge in $F$. Then this edge is in the graph $G\setminus F$. Using the triangle inequality, this edge is a shortest path between $a$ and $b$. Thus, the lemma holds.
Now assume that $\{a,b\}$ is in $F$. Let $c$ be a point in $C_{ab}$ as in Lemma 3. The path $P=(a,c,b)$ is in $G\setminus F$. Let $Q$ be a shortest path between $a$ and $b$ in $G\setminus F$, and assume that $Q$ contains a vertex $v$ that is not in $C_{ab}\cup\{a,b\}$; refer to Figure 3. Then
$|P|=|ac|+|cb|\leq|av|+|vb|\leq|Q|$
and, thus, $|P|=|Q|$. The path $P$ satisfies the lemma.
Figure 3: Illustrating the proof of Lemma 4.
The next lemma states that, in the graph $G\setminus F$, any two points $p$ and $q$ are connected by a “short” path, if their shortest path in $G^{\prime}$ contains “many” edges.
Lemma 5
Let $p$ and $q$ be two distinct points in $S$, and let $P^{\prime}$ be a shortest path between $p$ and $q$ in the graph $G^{\prime}$. If the number of vertices of $P^{\prime}$ is at least $2f+2$, then
$\delta_{G\setminus F}(p,q)\leq(8f+2)\cdot|P^{\prime}|.$
Proof. Let $e=\{a,b\}$ be any edge on the path $P^{\prime}$. By Lemma 4, there exists a shortest path $P_{e}$ between $a$ and $b$ in the graph $G\setminus F$, all of whose vertices belong to $C_{ab}\cup\{a,b\}$; refer to Figure 4.
Figure 4: The black path $P^{\prime}$ is a shortest path between $p$ and $q$ in $G^{\prime}$. For each edge $e=\{a,b\}$ on $P^{\prime}$, the green path $P_{e}$ is a shortest path between $a$ and $b$ in $G\setminus F$. The concatenation of all paths $P_{e}$ is a walk $P$ between $p$ and $q$ in $G\setminus F$.
Let $P$ be the concatenation of all paths $P_{e}$, where $e$ ranges over all edges of $P^{\prime}$. Then $P$ is a walk between $p$ and $q$ in $G\setminus F$. We will prove that the length of $P$ is at most $(8f+2)\cdot|P^{\prime}|$. This will prove the lemma.
Figure 5: $v_{e}$ is a vertex of $P^{\prime}$ that is closest (along $P^{\prime}$ and in terms of the number of edges) to $a$ or $b$, and that is not in $C_{ab}\cup\{a,b\}$.
Again, let $e=\{a,b\}$ be any edge on the path $P^{\prime}$. Recall that the set $C_{ab}\cup\{a,b\}$ has size $2f+1$ and $P^{\prime}$ contains at least $2f+2$ vertices. Therefore, there exists a vertex of $P^{\prime}$ that is not in $C_{ab}\cup\{a,b\}$. Let $v_{e}$ be such a vertex that is closest (along $P^{\prime}$ and in terms of the number of edges), to $a$ or $b$; refer to Figure 5. Then, again using the fact that the size of $C_{ab}\cup\{a,b\}$ is $2f+1$,
  • the subpath $P^{\prime}_{av_{e}}$ of $P^{\prime}$ between $a$ and $v_{e}$ has at most $2f+1$ edges, and
  • the subpath $P^{\prime}_{v_{e}b}$ of $P^{\prime}$ between $b$ and $v_{e}$ has at most $2f+1$ edges.
First assume that $e=\{a,b\}$ is an edge in $F$. Let $c$ be a point as in Lemma 3. Since both edges $\{a,c\}$ and $\{c,b\}$ are in $G\setminus F$, we have
$\delta_{G\setminus F}(a,b)\leq|ac|+|cb|.$
Since $v_{e}\not\in C_{ab}$, we have
$|ac|+|cb|\leq|av_{e}|+|v_{e}b|.$
By the triangle inequality, we have
$|av_{e}|+|v_{e}b|\leq|P^{\prime}_{av_{e}}|+|P^{\prime}_{v_{e}b}|.$
By combining the above three inequalities, we get
$\delta_{G\setminus F}(a,b)\leq|P^{\prime}_{av_{e}}|+|P^{\prime}_{v_{e}b}|.$ (7)
Observe that (7) also holds if $e=\{a,b\}$ is not an edge in $F$, because in that case, $\delta_{G\setminus F}(a,b)=|ab|$, which is at most equal to the right-hand side in (7).
Since $P$ is the concatenation of all paths $P_{e}$, with $e$ an edge of $P^{\prime}$, we have
$|P|$ $=$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}|P_{e}|$ (8)
$=$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}\delta_{G\setminus F}(a,b)$
$\leq$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}\left(|P^{\prime}_{av_{e}}|+|P^{\prime}_{v_{e}b}|\right)$
$=$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}|P^{\prime}_{av_{e}}|+\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}|P^{\prime}_{v_{e}b}|.$
Consider the first summation in (8). For each edge $\{x,y\}$ on the path $P^{\prime}$, let $N_{xy}$ be the number of edges $e=\{a,b\}$ on $P^{\prime}$ such that $\{x,y\}$ is an edge on the subpath $P^{\prime}_{av_{e}}$. Then
$\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}|P^{\prime}_{av_{e}}|=\sum_{\{x,y\}\mbox{\scriptsize{ on }}P^{\prime}}N_{xy}\cdot|xy|.$
To obtain an upper bound on $N_{xy}$, we assume, without loss of generality, that $x$ is on the subpath of $P^{\prime}$ between $p$ and $y$. Define the following subpath $Q_{xy}$ of $P^{\prime}$:
  • $\{x,y\}$ is an edge of $Q_{xy}$.
  • Starting at $x$, traverse $P^{\prime}$ towards $p$ and stop as soon as $2f$ edges have been traversed or the point $p$ has been reached. All edges traversed are edges of $Q_{xy}$.
  • Starting at $y$, traverse $P^{\prime}$ towards $q$ and stop as soon as $2f$ edges have been traversed or the point $q$ has been reached. All edges traversed are edges of $Q_{xy}$.
Recall that each subpath $P^{\prime}_{av_{e}}$ has at most $2f+1$ edges. It follows that every edge $e=\{a,b\}$ that is counted in $N_{xy}$ is an edge on $Q_{xy}$. Since $Q_{xy}$ has at most $4f+1$ edges, it follows that $N_{xy}\leq 4f+1$. Thus,
$\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}P^{\prime}}|P^{\prime}_{av_{e}}|\leq(4f+1)\sum_{\{x,y\}\mbox{\scriptsize{ on }}P^{\prime}}|xy|=(4f+1)\cdot|P^{\prime}|.$
By a symmetric argument, the second summation in (8) is at most $(4f+1)\cdot|P^{\prime}|$. We conclude that $|P|\leq(8f+2)|P^{\prime}|$.
We are now ready to complete the proof of Theorem 2. The proof of the next lemma uses the previous lemmas to obtain an upper bound on the length of a shortest path in $G\setminus F$ between any pairs of points.
Lemma 6
Let $p$ and $q$ be two distinct points in $S$. Then
$\delta_{G\setminus F}(p,q)\leq(8f+2)t\cdot\delta_{K_{S}\setminus F}(p,q).$
Proof. Let $Q$ be a shortest path between $p$ and $q$ in $K_{S}\setminus F$. By Lemma 2, $Q$ exists. For each edge $e=\{a,b\}$ on $Q$, let $P^{\prime}_{e}$ be a shortest path between $a$ and $b$ in $G^{\prime}$. Since $G^{\prime}$ is a $t$-spanner for $S$, we have $|P^{\prime}_{e}|\leq t|ab|$. We will show that
$\delta_{G\setminus F}(a,b)\leq(8f+2)\cdot|P^{\prime}_{e}|.$ (9)
This will imply that
$\delta_{G\setminus F}(p,q)$ $\leq$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}Q}\delta_{G\setminus F}(a,b)$
$\leq$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}Q}(8f+2)\cdot|P^{\prime}_{e}|$
$\leq$ $\sum_{e=\{a,b\}\mbox{\scriptsize{ on }}Q}(8f+2)t\cdot|ab|$
$=$ $(8f+2)t\cdot|Q|$
$=$ $(8f+2)t\cdot\delta_{K_{S}\setminus F}(p,q).$
If $P^{\prime}_{e}$ has at least $2f+2$ vertices, then (9) follows from Lemma 5. From now on, we assume that $P^{\prime}_{e}$ has at most $2f+1$ vertices and, thus, at most $2f$ edges.
Since $Q$ is a path in $K_{S}\setminus F$ and $e=\{a,b\}$ is an edge on $Q$, this edge is not in $F$. If $e$ is an edge in $G\setminus F$, then
$\delta_{G\setminus F}(a,b)=|ab|\leq|P^{\prime}_{e}|$
and, thus (9) holds.
It remains to consider the case when $e=\{a,b\}$ is not an edge in $G\setminus F$. Consider the edge $\{a,d\}$ on $P^{\prime}_{e}$ that is incident on $a$. Observe that $d\neq b$, because $\{a,b\}$ is not an edge in $G^{\prime}$.
We first show that for every point $z$ in $C_{ad}$,
$|az|+|zd|\leq|ab|+|bd|.$ (10)
The proof is by contradiction: If $|ab|+|bd|<|az|+|zd|$, then $b$ is an element of $C_{ad}$ and, therefore, by the definition of the graph $G$, $\{a,b\}$ is an edge in $G$, which is a contradiction.
Next, we show that
$\delta_{G\setminus F}(a,d)\leq 2\cdot|P^{\prime}_{e}|.$ (11)
If $\{a,d\}$ is not in $F$, then $\delta_{G\setminus F}(a,d)=|ad|\leq|P^{\prime}_{e}|$ and, thus, (11) holds. Assume that $\{a,d\}$ is an edge in $F$. By Lemma 3, there exists a point $c$ in $C_{ad}$, such that both $\{a,c\}$ and $\{c,d\}$ are edges in $G\setminus F$. Then, $\delta_{G\setminus F}(a,d)\leq|ac|+|cd|$, which, by (10), is at most $|ab|+|bd|$, which is at most $2\cdot|P^{\prime}_{e}|$, i.e., (11) holds.
Let $D=C_{ad}\cup\{a,d\}$. Observe that $D$ has size $2f+1$. Let $\{x,y\}$ be an arbitrary edge on $P^{\prime}_{e}$ with $\{x,y\}\neq\{a,d\}$. Let $D^{\prime}=D\setminus\{x,y\}$. Then the size of $D^{\prime}$ is at least 2f12f-1. Let $d^{\prime}$ be a point in $D^{\prime}$ for which $|xd^{\prime}|+|d^{\prime}y|$ is maximum. Refer to Figure 6.
Figure 6: $D=C_{ad}\cup\{a,d\}$, $\{x,y\}$ is an edge on $P^{\prime}_{e}$, $\{x,y\}\neq\{a,d\}$.
First, we assume that $\{x,y\}$ is an edge in $F$. By Lemma 3, there is a point $c^{\prime}$ in $C_{xy}$ such that both $\{x,c^{\prime}\}$ and $\{c^{\prime},y\}$ are edges in $G\setminus F$. It is clear that
$\delta_{G\setminus F}(x,y)\leq|xc^{\prime}|+|c^{\prime}y|.$
We next argue that
$|xc^{\prime}|+|c^{\prime}y|\leq|xd^{\prime}|+|d^{\prime}y|.$
Assume, by contradiction, that $|xd^{\prime}|+|d^{\prime}y|<|xc^{\prime}|+|c^{\prime}y|$. Then for all $z$ in $D^{\prime}=D\setminus\{x,y\}$, $|xz|+|zy|<|xc^{\prime}|+|c^{\prime}y|$. There are at least 2f12f-1 such points $z$, and each of them is in $S\setminus\{x,y\}$. Thus, $c^{\prime}\not\in C_{xy}$, which is a contradiction.
The previous two inequalities, together with the triangle inequality, imply that
$\delta_{G\setminus F}(x,y)\leq|xd^{\prime}|+|d^{\prime}y|$ $\leq$ $\left(|P^{\prime}_{xa}|+|ad^{\prime}|\right)+\left(|d^{\prime}d|+|P^{\prime}_{dy}|\right)$
$\leq$ $|ad^{\prime}|+|d^{\prime}d|+2\cdot|P^{\prime}_{e}|.$
We claim that
$|ad^{\prime}|+|d^{\prime}d|\leq|ab|+|bd|.$ (12)
If $d^{\prime}\in\{a,d\}$, then (12) follows from the triangle inequality. If $d^{\prime}\not\in\{a,d\}$, then $d^{\prime}\in C_{ad}$ and (12) follows from (10).
Since $|ab|+|bd|\leq 2\cdot|P^{\prime}_{e}|$, we conclude that
$\delta_{G\setminus F}(x,y)\leq 4\cdot|P^{\prime}_{e}|.$ (13)
In the derivation of (13), we assumed that $\{x,y\}$ is an edge in $F$. If $\{x,y\}$ is not an edge in $F$, then (13) also holds, because, in that case, $\delta_{G\setminus F}(x,y)=|xy|\leq|P^{\prime}_{e}|$.
For each edge $e^{\prime}=\{x,y\}$ on $P^{\prime}_{e}$, let $P_{e^{\prime}}$ be a shortest path between $x$ and $y$ in $G\setminus F$. Let $P_{ab}$ be the concatenation of all paths $P_{e^{\prime}}$, where $e^{\prime}$ ranges over all edges of $P^{\prime}_{e}$. Observe that $P_{ab}$ is a walk between $a$ and $b$ in $G\setminus F$. We have
$\delta_{G\setminus F}(a,b)$ $\leq$ $|P_{ab}|$
[…]
$=$ $\sum_{e^{\prime}=\{x,y\}\mbox{\scriptsize{ on }}P^{\prime}_{e}}\delta_{G\setminus F}(x,y)$
$\leq$ $\sum_{e^{\prime}=\{x,y\}\mbox{\scriptsize{ on }}P^{\prime}_{e}}4\cdot|P^{\prime}_{e}|.$
Since $P^{\prime}_{e}$ has at most $2f$ edges, the latter summation has at most $2f$ terms. We conclude that
$\delta_{G\setminus F}(a,b)\leq 8f\cdot|P^{\prime}_{e}|$
and, thus, (9) holds.
2.2.1 A Lower Bound for our Construction
We have shown that any $t$-spanner $G^{\prime}$ can be transformed to an $f$-faulty-degree $t^{\prime}$-spanner with $t^{\prime}=O(ft)$. Below, we will give an example for which $t^{\prime}=\Omega(ft)$.
Let $f\geq 3$ be an integer and let $\varepsilon$ be a real number with $0<\varepsilon\ll 1/(f-2)$. Let $B_{0},B_{1},\ldots B_{f}$ be pairwise disjoint sets, where $B_{0}=\{b_{0},b_{1},\ldots,b_{f}\}$, and each $B_{i}$, for $1\leq i\leq f$, consists of $2f$ points. Let $S$ be the union of $B_{0},B_{1},\ldots B_{f}$.
We define the following graph $H$ with vertex set $S$; refer to Figure 7:
Figure 7: Illustrating the graph $H$. All black edges have weight 11, all blue edges have weight $\varepsilon$, and the dotted edge $\{b_{0},b_{f}\}$ has weight $1+\varepsilon$. The set $B_{0}$ is equal to $\{b_{0},b_{1},\ldots,b_{f}\}$. The graph $G^{\prime}$ consists of all solid edges.
  • $H$ contains the edge $\{b_{0},b_{f}\}$ having weight $1+\varepsilon$.
  • $H$ contains the edges $\{b_{0},b_{1}\}$ and $\{b_{f-1},b_{f}\}$, each having weight $1$.
  • $H$ contains the edges $\{b_{j},b_{j+1}\}$, for $j=1,2,\ldots,f-2$, each having weight $\varepsilon$.
  • For each $i$ with $1\leq i\leq f$, $H$ contains the complete graph on $B_{i}$, with each edge having weight $\varepsilon$.
  • For each $i$ with $1\leq i\leq f$, $H$ contains the complete bipartite graph on $B_{i}\times\{b_{i-1},b_{i}\}$, with each edge having weight $1$.
For any two points $p$ and $q$ of $S$, we define $|pq|$ to be the shortest-path distance between $p$ and $q$ in the graph $H$. It is clear that $(S,|\cdot|)$ is a metric space.
Observation 1
For each edge $\{p,q\}$ in $H$, $|pq|$ is the weight of this edge, i.e., $\{p,q\}$ is the shortest path between $p$ and $q$ in $H$.
Proof. The proof can be obtained by verifying that for each edge $\{p,q\}$, any path between $p$ and $q$ that has at least two edges has length more than the weight of $\{p,q\}$.
Let $G^{\prime}$ be the graph obtained from $H$ by removing the edge $\{b_{0},b_{f}\}$.
Observation 2
$G^{\prime}$ is a $3$-spanner for $S$.
Proof. If $\{x,y\}$ is an edge of $H$ and $\{x,y\}\neq\{b_{0},b_{f}\}$, then, by Observation 1, $\delta_{G^{\prime}}(x,y)=|xy|$. For the edge $\{b_{0},b_{f}\}$, we have, using the fact that $\varepsilon<1/(f-2)$,
$\delta_{G^{\prime}}(b_{0},b_{f})=1+(f-2)\varepsilon+1\leq 3\leq 3\cdot|b_{0}b_{f}|.$
It follows that, for any two points $p$ and $q$ in $S$,
$\delta_{G^{\prime}}(p,q)\leq 3\cdot\delta_{H}(p,q)=3\cdot|pq|.$
Having defined the $3$-spanner $G^{\prime}$, consider the graph $G$ obtained from $G^{\prime}$ by applying the transformation given in the beginning of Section 2.2.
Observation 3
$\{b_{0},b_{f}\}$ is not an edge in the graph $G$.
Proof. Let $\{p,q\}$ be any edge in $G^{\prime}$. Recall that $G$ contains the complete graph with vertex set $C_{pq}\cup\{p,q\}$. The claim holds if we can show that at least one of $b_{0}$ and $b_{f}$ is not in $C_{pq}\cup\{p,q\}$.
Observe that $|pb_{0}|+|b_{0}q|\geq 2+\varepsilon$ or $|pb_{f}|+|b_{f}q|\geq 2+\varepsilon$, while there are at least $2f$ points $x$ in $S\setminus\{p,q\}$ such that $|px|+|xq|\leq 2$. This implies that at least one of $b_{0}$ and $b_{f}$ is not in $C_{pq}\cup\{p,q\}$.
Observation 4
For any $i$ and $j$ with $1\leq i<j\leq f$, the graph $G$ does not contain any edge between the sets $B_{i}$ and $B_{j}$.
Proof. For any point $a$ in $B_{i}$ and any point $b$ in $B_{j}$, we have $|ab|\geq 2$. For any edge $\{p,q\}$ in $G^{\prime}$, we have $|pa|+|aq|\geq 2+\varepsilon$ or $|pb|+|bq|\geq 2+\varepsilon$, while there are at least $2f$ points $x$ in $S\setminus\{p,q\}$ such that $|px|+|xq|\leq 2$. This implies that at least one of $a$ and $b$ is not in $C_{pq}\cup\{p,q\}$.
By Theorem 2, the graph $G$ is an $f$-faulty-degree $t^{\prime}$-spanner for $S$, where $t^{\prime}\leq 3(8f+2)=24f+6$. Below, we show that $t^{\prime}\geq f$.
Let $F$ be the set of all edges $\{p,q\}$ of $G$ for which both $p$ and $q$ are in $B_{0}$. Since $B_{0}$ has size $f+1$, the maximum degree of the graph $(S,F)$ is at most $f$.
By Observation 3, $\{b_{0},b_{f}\}$ is not an edge in $G\setminus F$. This, together with Observation 4, implies that the shortest path between $b_{0}$ and $b_{f}$ in $G\setminus F$ consists of the following:
  • For $i=0,1,\ldots,f-1$, it takes the edge from $b_{i}$ to an arbitrary point, say $x_{i+1}$, of $B_{i+1}$, and then the edge from $x_{i+1}$ to $b_{i+1}$.
Thus,
$\delta_{G\setminus F}(b_{0},b_{f})=2f.$
Since by Observation 3, $\{b_{0},b_{f}\}$ is not in $F$, we have
$\delta_{K_{S}\setminus F}(b_{0},b_{f})=|b_{0}b_{f}|=1+\varepsilon\leq 2.$
It follows that
$\frac{\delta_{G\setminus F}(b_{0},b_{f})}{\delta_{K_{S}\setminus F}(b_{0},b_{f})}\geq f.$
2.2.2 Comparison with the Bodwin–Haeupler–Parter Construction
Recall that Bodwin, Haeupler, and Parter [4] proved the following: Let $H$ be any edge-weighted graph with $n$ vertices. For any positive integers $f$ and $k$, there exists an $f$-faulty-degree (2k1)(2k-1)-spanner for $H$ that has
$O\left(k^{k}\cdot f^{1-1/k}\cdot n^{1+1/k}\right)$
edges. In particular, this results holds when $H$ is the complete graph induced by a metric space with $n$ points.
Let $S$ be an arbitrary metric space with $n$ points. For any integer $k^{\prime}\geq 1$, the greedy algorithm of Althöfer et al. [2] constructs a $(2k^{\prime}-1)$-spanner for $S$ with $O(n^{1+1/k^{\prime}})$ eges. By taking $k^{\prime}=\Theta(k/f)$ and applying Theorem 2 to this spanner, we obtain an $f$-faulty-degree $O(k)$-spanner for $S$, whose number of edges is
$O\left(f\cdot n^{1+1/k^{\prime}}\right)=n^{\Theta(f/k)}\cdot f\cdot n^{1+1/k}.$
If $k=\omega(\sqrt{f\cdot\log n})$, then $n^{\Theta(f/k)}=o(k^{k})$ and, thus, the number of edges in our construction is smaller than the number of edges in the Bodwin–Haeupler–Parter construction.
3 Metric Spaces with a Small WSPD
The construction in the previous section is valid for any metric space that admits a $t$-spanner, for some value of $t$. However, the stretch factor of the $f$-faulty-degree spanner is $\Theta(ft)$. In this section, we will show that, if the metric space admits a well-separated pair decomposition, we can obtain a stretch factor of $1+\varepsilon$, for any given $\varepsilon>0$.
3.1 Preliminaries
Let $(S,|\cdot|)$ be a finite metric space. For any two non-empty subsets $A$ and $B$ of $S$, we define their distance $|AB|$ as
$|AB|=\min\{|pq|:p\in A,q\in B\}$
and the diameter $\mathord{\it diam}(A)$ of $A$ as
$\mathord{\it diam}(A)=\max\{|pq|:p\in A,q\in A\}.$
For a real number $c>0$, called the separation ratio, we say that $A$ and $B$ are well-separated, if
$|AB|\geq c\cdot\max(\mathord{\it diam}(A),\mathord{\it diam}(B)).$
Lemma 7
Let $c>0$ be a real number, let $A$ and $B$ be two non-empty subsets of $S$ that are well-separated with respect to $c$, let $p$ and $p^{\prime}$ be any two points in $A$, and let $q$ and $q^{\prime}$ be any two points in $B$. Then
  1. 1.
    $|pp^{\prime}|\leq(1/c)|pq|$,
  2. 2.
    $|qq^{\prime}|\leq(1/c)|pq|$, and
  3. 3.
    $|p^{\prime}q^{\prime}|\leq(1+2/c)|pq|$.
Proof. The first claim follows from the following chain of inequalities:
$|pp^{\prime}|$ $\leq$ $\mathord{\it diam}(A)$
$\leq$ $\max(\mathord{\it diam}(A),\mathord{\it diam}(B))$
$\leq$ $(1/c)|AB|$
$\leq$ $(1/c)|pq|.$
The proof of the second claim is symmetric. For the proof of the third claim, we combine the first two claims and use the triangle inequality:
$|p^{\prime}q^{\prime}|$ $\leq$ $|p^{\prime}p|+|pq|+|qq^{\prime}|$
$\leq$ $(1/c)|pq|+|pq|+(1/c)|pq|$
$=$ $(1+2/c)|pq|.$
Definition 2 (Callahan and Kosaraju [5])
Let $(S,|\cdot|)$ be a finite metric space and let $c>0$ be a real number. A well-separated pair decomposition (WSPD) for $S$ is a sequence
$\{A_{1},B_{1}\},\{A_{2},B_{2}\},\ldots,\{A_{m},B_{m}\}$
of pairs of non-empty subsets of $S$, for some integer $m$, such that
  1. 1.
    for each $i$ with $1\leq i\leq m$, $A_{i}$ and $B_{i}$ are well-separated, and
  2. 2.
    for any two distinct points $p$ and $q$ of $S$, there is exactly one index $i$ such that
    1. (a)
      $p\in A_{i}$ and $q\in B_{i}$, or
    2. (b)
      $p\in B_{i}$ and $q\in A_{i}$.
The integer $m$ is called the size of the well-separated pair decomposition.
3.2 The Faulty-Degree Spanner
Let $(S,|\cdot|)$ be a finite metric space, let $f\geq 1$ be an integer, and let $\varepsilon>0$ be a real number. Assume that
$\{A_{1},B_{1}\},\{A_{2},B_{2}\},\ldots,\{A_{m},B_{m}\}$
is a well-separated pair decomposition for $S$ with separation ratio
$c=2+4/\varepsilon.$
We define a graph $G=(S,E)$ whose edge set $E$ is obtained as follows: For each $i$ with $1\leq i\leq m$, define the following two subsets $A^{\prime}_{i}$ and $B^{\prime}_{i}$ of $A_{i}$ and $B_{i}$, respectively:
  • If $A_{i}$ has size at most $2f+1$, then $A^{\prime}_{i}=A_{i}$.
  • If $A_{i}$ has size at least $2f+2$, then $A^{\prime}_{i}$ is an arbitrary subset of size $2f+1$ of $A_{i}$.
  • If $B_{i}$ has size at most $2f+1$, then $B^{\prime}_{i}=B_{i}$.
  • If $B_{i}$ has size at least $2f+2$, then $B^{\prime}_{i}$ is an arbitrary subset of size $2f+1$ of $B_{i}$.
The edge set $E$ contains the edges of the complete bipartite graph with vertex sets $A^{\prime}_{i}$ and $B^{\prime}_{i}$.
It is clear that the total number of edges in $E$ is at most $(2f+1)^{2}m$. In the following subsection, we will prove that $G$ is an $f$-faulty-degree $(1+\varepsilon)$-spanner for $S$.
3.3 Analysis
Let $F$ be an arbitrary subset of the edge set of $G$ such that the maximum degree of the graph $(S,F)$ is at most $f$.
Lemma 8
For each edge $\{p,q\}$ in $K_{S}\setminus F$, we have
$\delta_{G\setminus F}(p,q)\leq(1+\varepsilon)\cdot|pq|.$
Proof. Let $i$ be the index in $\{1,2,\ldots,m\}$ such that (a) $p\in A_{i}$ and $q\in B_{i}$ or (b) $p\in B_{i}$ and $q\in A_{i}$. We may assume without loss of generality that (a) holds.
The proof is by induction on the rank of the distance $|pq|$ in the set of all edge lengths in $K_{S}\setminus F$.
For the base case, assume that $\{p,q\}$ is a shortest edge in $K_{S}\setminus F$. We first show that the size of the set $A_{i}$ is at most $2f$. Assume that this set has at least $2f+1$ points. Then $K_{S}$ contains at least $2f\geq f+1$ edges $\{p,p^{\prime}\}$ for which $p^{\prime}\in A_{i}$. At least one of these edges is in the graph $K_{S}\setminus F$. Since by Lemma 7, $|pp^{\prime}|<|pq|$, this is a contradiction. By a symmetric argument, we can show that the size of the set $B_{i}$ is at most $2f$. It follows that $p\in A_{i}=A^{\prime}_{i}$ and $q\in B_{i}=B^{\prime}_{i}$. Thus, $\{p,q\}$ is an edge in $G$. Since $\{p,q\}\not\in F$, it follows that
$\delta_{G\setminus F}(p,q)=|pq|\leq(1+\varepsilon)\cdot|pq|.$
For the induction step, assume that $\{p,q\}$ is not a shortest edge in $K_{S}\setminus F$. If $\{p,q\}$ is in $G$, then it is in $G\setminus F$ and the lemma holds. Assume from now on that $\{p,q\}$ is not an edge in $G$. Then $p\not\in A^{\prime}_{i}$ or $q\not\in B^{\prime}_{i}$. We distinguish three cases.
Case 1: $p\not\in A^{\prime}_{i}$ and $q\not\in B^{\prime}_{i}$.
We observe that each of $A_{i}$ and $B_{i}$ has at least $2f+2$ points, and each of $A^{\prime}_{i}$ and $B^{\prime}_{i}$ has exactly $2f+1$ points.
Since $K_{S}$ contains $2f+1\geq f+1$ edges $\{p,p^{\prime}\}$ for which $p^{\prime}$ is in $A^{\prime}_{i}$, there is a point $p^{\prime}$ in $A^{\prime}_{i}$ such that $\{p,p^{\prime}\}$ is in $K_{S}\setminus F$. Since, by Lemma 7, $|pp^{\prime}|\leq(1/c)|pq|<|pq|$, we have, by induction,
$\delta_{G\setminus F}(p,p^{\prime})\leq(1+\varepsilon)\cdot|pp^{\prime}|.$
Note that, by the definition of $G$, for each point $q^{\prime}$ in $B^{\prime}_{i}$, the edge $\{p^{\prime},q^{\prime}\}$ is in $G$. Define
$B^{\prime\prime}_{i}=\{q^{\prime}\in B^{\prime}_{i}:\{p^{\prime},q^{\prime}\}\mbox{ is an edge in }G\setminus F\}.$
Since the size of $B^{\prime}_{i}$ is $2f+1$, the size of $B^{\prime\prime}_{i}$ is at least $f+1$. Thus, $K_{S}$ contains at least $f+1$ edges $\{q,q^{\prime}\}$ for which $q^{\prime}$ is in $B^{\prime\prime}_{i}$. It follows that there is a point $q^{\prime}$ in $B^{\prime\prime}_{i}$ such that $\{q,q^{\prime}\}$ is in $K_{S}\setminus F$. Since, by Lemma 7, $|qq^{\prime}|\leq(1/c)|pq|<|pq|$, we have, by induction,
$\delta_{G\setminus F}(q,q^{\prime})\leq(1+\varepsilon)\cdot|qq^{\prime}|.$
Note that, since $q^{\prime}$ is in $B^{\prime\prime}_{i}$, $\{p^{\prime},q^{\prime}\}$ is an edge in $G\setminus F$. It follows that
$\delta_{G\setminus F}(p,q)$ $\leq$ $\delta_{G\setminus F}(p,p^{\prime})+|p^{\prime}q^{\prime}|+\delta_{G\setminus F}(q^{\prime},q)$
$\leq$ $(1+\varepsilon)\cdot|pp^{\prime}|+|p^{\prime}q^{\prime}|+(1+\varepsilon)\cdot|q^{\prime}q|.$
Using Lemma 7, we conclude that
$\delta_{G\setminus F}(p,q)$ $\leq$ $(1+\varepsilon)\cdot(1/c)|pq|+(1+2/c)|pq|+(1+\varepsilon)\cdot(1/c)|pq|$
$=$ $(1+(4+2\varepsilon)/c)|pq|$
$=$ $(1+\varepsilon)|pq|,$
where the last equality follows from our choice for $c$.
Case 2: $p\in A^{\prime}_{i}$ and $q\not\in B^{\prime}_{i}$.
In this case, $B_{i}$ has at least $2f+2$ points and $B^{\prime}_{i}$ has exactly $2f+1$ points. By the definition of $G$, for each point $q^{\prime}$ in $B^{\prime}_{i}$, the edge $\{p,q^{\prime}\}$ is in $G$. Define
$B^{\prime\prime}_{i}=\{q^{\prime}\in B^{\prime}_{i}:\{p,q^{\prime}\}\mbox{ is an edge in }G\setminus F\}.$
Note that the size of $B^{\prime\prime}_{i}$ is at least $f+1$. Thus, $K_{S}$ contains at least $f+1$ edges $\{q,q^{\prime}\}$ for which $q^{\prime}$ is in $B^{\prime\prime}_{i}$. It follows that there is a point $q^{\prime}$ in $B^{\prime\prime}_{i}$ such that $\{q,q^{\prime}\}$ is in $K_{S}\setminus F$. By Lemma 7, $|qq^{\prime}|<|pq|$, implying that
$\delta_{G\setminus F}(q^{\prime},q)\leq(1+\varepsilon)\cdot|q^{\prime}q|.$
Since $q^{\prime}$ is in $B^{\prime\prime}_{i}$, $\{p,q^{\prime}\}$ is an edge in $G\setminus F$. It follows that
$\delta_{G\setminus F}(p,q)\leq|pq^{\prime}|+\delta_{G\setminus F}(q^{\prime},q)\leq|pq^{\prime}|+(1+\varepsilon)\cdot|q^{\prime}q|.$
As in Case 1, it follows that
$\delta_{G\setminus F}(p,q)\leq(1+\varepsilon)\cdot|pq|.$
Case 3: $p\not\in A^{\prime}_{i}$ and $q\in B^{\prime}_{i}$.
This case is symmetric to Case 2.
We summarize the result of this section in the following theorem:
Theorem 3
Let $(S,|\cdot|)$ be a finite metric space, let $f\geq 1$ be an integer, and let $\varepsilon>0$ be a real number. Assume there exists a well-separated pair decomposition for $S$ with separation ratio $c=2+4/\varepsilon$. Let $m$ be the size of this decomposition. Then the graph $G=(S,E)$ defined at the beginning of Section 3.2 is an $f$-faulty-degree $(1+\varepsilon)$-spanner for $S$ that has at most $(2f+1)^{2}m$ edges.
Proof. Let $F$ be an arbitrary subset of the edge set of $G$ such that the maximum degree of the graph $(S,F)$ is at most $f$. Let $p$ and $q$ be two distinct points in $S$. To prove that
$\delta_{G\setminus F}(p,q)\leq(1+\varepsilon)\cdot\delta_{K_{S}\setminus F}(p,q),$
we consider a shortest path in $K_{S}\setminus F$ and apply Lemma 8 to each edge on this path.
We present some examples for which Theorem 3 implies non-trivial results. Callahan and Kosaraju [5] have shown that for any set $S$ of $n$ points in $\mathbb{R}^{d}$, where the dimension $d$ is a constant and $|\cdot|$ is the Euclidean distance function, there is a WSPD for $S$ of size $m=O(n)$. Har-Peled and Mendel [12] generalized this this by showing that for every metric space of bounded doubling dimension, there is a WSPD for $S$ of size $m=O(n)$. Theorem 3 implies that in both of these cases, we can construct an $f$-faulty-degree $(1+\varepsilon)$-spanner with $O(f^{2}n)$ edges. In the next section, we will improve the number of edges to $O(fn)$ for the Euclidean case.
Let $S$ be a set of $n$ points in $\mathbb{R}^{d}$, where $d\geq 2$ is a constant. The unit-disk graph of $S$ is the graph with vertex set $S$, in which any two distinct points $p$ and $q$ are connected by an edge if their Euclidean distance $|pq|$ is at most one; the weight of this edge is equal to $|pq|$. The shortest-path distance in this graph defines a metric space. Note that the doubling dimension of this space is not bounded. Gao and Zhang [11] have proved the following results:
  • For the case when $d=2$, there is a WSPD for the unit-disk graph metric on $S$ of size $m=O(n\log n)$. Theorem 3 implies that there exists an $f$-faulty-degree $(1+\varepsilon)$-spanner with $O(f^{2}n\log n)$ edges.
  • For the case when $d\geq 3$, there is a WSPD for the unit-disk graph metric on $S$ of size $m=O(n^{2-2/d})$. Theorem 3 implies that there exists an $f$-faulty-degree $(1+\varepsilon)$-spanner with $O(f^{2}n^{2-2/d})$ edges.
4 Euclidean Metric Spaces of Constant Dimension
In this section, we consider the metric space $(S,|\cdot|)$, where $S$ is a set of $n$ points in $\mathbb{R}^{d}$, for some constant $d$, and $|\cdot|$ is the Euclidean distance function.
We will define two graphs with vertex set $S$ that are variants of the Yao-graph (see [10, 21]) and the $\Theta$-graph (see [8, 14]) and that depend on the maximum degree bound $f$ of any set $F$ of faulty edges. As we will see, for any real constant $\varepsilon>0$, each of these graphs is an $f$-faulty-degree $(1+\varepsilon)$-spanner with $O(fn)$ edges, where $n$ is the size of $S$.
4.1 Cones
Let $\theta$ be a real number with $0<\theta<\pi/4$ and let $\mathcal{C}$ be a collection of cones that cover $\mathbb{R}^{d}$, such that for each cone $C$ in $\mathcal{C}$, the apex of $C$ is at the origin 0 and the angular diameter of $C$ is at most $\theta$, i.e.,
$\max\left\{\angle(0x,0y):x,y\in C\setminus\{0\}\right\}\leq\theta.$
Lukovszki [18] has shown how to obtain such a collection $\mathcal{C}$ consisting of $O(1/\theta^{d-1})$ cones; see also [19, Chapter 5].
4.2 Yao- and Theta-Graphs
Let $S$ be a set of $n$ points in $\mathbb{R}^{d}$ and let $f\geq 0$ be an integer. Let $\theta$ be a real constant with $0<\theta<\pi/4$, and let $\mathcal{C}$ be a collection of cones as in Section 4.1.
For any cone $C$ in $\mathcal{C}$, we fix an arbitrary ray $\ell_{C}$ that emanates from the origin and is completely contained in $C$. For any point $p$ in $S$ and any cone $C$ in $\mathcal{C}$, let $C+p$ denote the cone with apex $p$ obtained by translating $C$, i.e., $C+p=\{x+p:x\in C\}$. Similarly, we denote by $\ell_{C}+p$ the ray emanating from $p$ that is parallel to $\ell_{C}$. We denote by $S_{p,C}$ the set of all points in $S\setminus\{p\}$ that are in the cone $C+p$, i.e., $S_{p,C}=(C+p)\cap(S\setminus\{p\})$.
We are now ready to define the variants of the Yao- and $\Theta$-graphs; refer to Figure 8.
[…]
  1. 1.
[…]
  2. 2.
    The graph $\Theta(\theta,2f+1)$ also has vertex set $S$. For each point $p$ in $S$ and for each cone $C$ in $\mathcal{C}$, consider again the set $S_{p,C}$. We add an edge between $p$ and each of the $\min(2f+1,|S_{p,C}|)$ points in $S_{p,C}$, whose orthogonal projections onto the ray $\ell_{C}+p$ are closest (with respect to the Euclidean distance) to $p$.
It is clear that the number of edges in each of these graphs is at most $|\mathcal{C}|\cdot(2f+1)n=O(fn)$.
4.3 Analysis
[…]
We remark that Lukovszki [17] used $\Theta(\theta,f+1)$ for the case when at most $f$ edges can be removed. Our proof is similar.
Our proof will use the following geometric result, which is illustrated in Figure 9.
Figure 9: Illustrating the first case in Lemma 9.
Lemma 9
Let $p$ and $q$ be two distinct points in $S$, and let $C$ be a cone in $\mathcal{C}$ such that $q\in C+p$.
  1. 1.
    Let $\ell$ be an arbitrary ray that emanates from $p$ and is completely contained in $C+p$. Let $r$ be a point in $S_{p,C}$ such that the orthogonal projection of $r$ onto $\ell$ is at least as close to $p$ as the orthogonal projection of $q$ onto $\ell$. Then,
    $|rq|\leq|pq|-(\cos\theta-\sin\theta)|pr|.$
  2. 2.
    Let $r$ be a point in $S_{p,C}$ such that $|pr|\leq|pq|$. Then,
    $|rq|\leq|pq|-(\cos\theta-\sin\theta)|pr|.$
Proof. The proof of the first claim can be found in [19, Lemma 4.1.4]. The second claim follows from the first claim, by taking for $\ell$ the ray that emanates from $p$ and goes through $q$.
[…]
For simplicity, we denote the graph $\Theta(\theta,2f+1)$ by $\Theta$. Let $F$ be an arbitrary subset of the edge set of $\Theta$ and assume that the maximum degree of the graph $(S,F)$ is at most $f$.
Lemma 10
Let $\{p,q\}$ be an edge in $K_{S}\setminus F$ that is not in $\Theta$. Then, there is a point $r$ in $S_{p,C}$ such that
  • $r\neq q$,
  • the orthogonal projection of $r$ onto the ray $\ell_{C}+p$ is at least as close to $p$ as the orthogonal projection of $q$ onto $\ell_{C}+p$,
  • $\{p,r\}$ is an edge in $\Theta\setminus F$, and
  • $\{r,q\}$ is an edge in $K_{S}\setminus F$.
Proof. Since $\{p,q\}$ is not an edge in $\Theta$, this graph contains $2f+1$ edges $\{p,r\}$, where $r$ is in $S_{p,C}$ and $r\neq q$. For each such edge $\{p,r\}$, the orthogonal projection of $r$ onto the ray $\ell_{C}+p$ is at least as close to $p$ as the orthogonal projection of $q$ onto $\ell_{C}+p$. Since the degree of $p$ in $F$ is at most $f$, at least $f+1$ of these edges $\{p,r\}$ are in $\Theta\setminus F$. Observe that for each such edge $\{p,r\}$, the edge $\{r,q\}$ is in $K_{S}$. Since the degree of $q$ in $F$ is at most $f$, for at least one of these $f+1$ edges $\{p,r\}$ that are in $\Theta\setminus F$, the edge $\{r,q\}$ is in $K_{S}\setminus F$.
Lemma 11
For every edge $\{p,q\}$ in $K_{s}\setminus F$, $\delta_{\Theta\setminus F}(p,q)\leq t\cdot|pq|$, where $t=1/(\cos\theta-\sin\theta)$.
Proof. Let $C$ be a cone in $\mathcal{C}$ such that $q\in C+p$. We will prove the lemma by induction on the rank of the Euclidean distance $|pq|$ in the set of all edge lengths in $K_{S}\setminus F$.
For the base case, assume that $\{p,q\}$ is a shortest edge in $K_{s}\setminus F$. We claim that this edge is in $\Theta$. If this is true, then it is in $\Theta\setminus F$ as well and $\delta_{\Theta\setminus F}(p,q)=|pq|\leq t\cdot|pq|$. Assume that $\{p,q\}$ is not in $\Theta$. Let $r$ be a point in $S_{p,C}$ as in Lemma 10. It follows from the first claim in Lemma 9, with $\ell=\ell_{C}+p$, that $|rq|<|pq|$, which contradicts the assumption that $\{p,q\}$ is a shortest edge in $K_{s}\setminus F$.
From now on, assume that $\{p,q\}$ is not a shortest edge in $K_{s}\setminus F$. If this edge is in $\Theta$, then it is in $\Theta\setminus F$ and the lemma holds. Otherwise, let $r$ be a point in $S_{p,C}$ as in Lemma 10. As above, by Lemma 9, we have $|rq|<|pq|$. Thus, by induction,
$\delta_{\Theta\setminus F}(r,q)\leq t\cdot|rq|.$
It follows that
$\delta_{\Theta\setminus F}(p,q)\leq|pr|+\delta_{\Theta\setminus F}(r,q)\leq|pr|+t\cdot|rq|.$
Using the first claim in Lemma 9, with $\ell=\ell_{C}+p$, a straightforward calculation shows that
$\delta_{\Theta\setminus F}(p,q)\leq t\cdot|pq|.$
Theorem 4
[…]
Proof. We only give the proof for $\Theta=\Theta(\theta,2f+1)$. Set $t=1/(\cos\theta-\sin\theta)$. Let $F$ be an arbitrary subset of the edge set of $\Theta$, such that the maximum degree of the graph $(S,F)$ is at most $f$. Let $p$ and $q$ be two distinct points in $S$. If $\{p,q\}$ is an edge in $K_{S}\setminus F$, then, by Lemma 11,
$\delta_{\Theta\setminus F}(p,q)\leq t\cdot|pq|=t\cdot\delta_{K_{s}\setminus F}(p,q)\leq(1+\varepsilon)\cdot\delta_{K_{s}\setminus F}(p,q).$
If $\{p,q\}$ is not an edge in $K_{S}\setminus F$, then we consider a shortest path between $p$ and $q$ in $K_{S}\setminus F$, and apply Lemma 11 to each edge on this path.
Acknowledgement
We thank the anonymous referees for their careful reading of the first version of this paper, and for their detailed comments and suggestions.